### v1

# Depth-Width Tradeoffs in Algorithmic Reasoning of Graph Tasks with Transformers

Abstract
Transformers have revolutionized the field of machine learning. In particular, they can be used to solve complex algorithmic problems, including graph-based tasks. In such algorithmic tasks a key question is what is the minimal size of a transformer that can implement a task. Recent work has begun to explore this problem for graph-based tasks, showing that for sub-linear embedding dimension (i.e., model width) logarithmic depth suffices. However, an open question, which we address here, is what happens if width is allowed to grow linearly. Here we analyze this setting, and provide the surprising result that with linear width, constant depth suffices for solving a host of graph-based problems. This suggests that a moderate increase in width can allow much shallower models, which are advantageous in terms of inference time. For other problems, we show that quadratic width is required. Our results demonstrate the complex and intriguing landscape of transformer implementations of graph-based algorithms. We support our theoretical results with empirical evaluations.
Machine Learning, ICML
1 Introduction
The transformer architecture (Vaswani, 2017), which was initially introduced for machine translation (Bahdanau, 2014), has emerged as the state-of-the-art neural network architecture across many fields, including computer vision (Dosovitskiy et al., 2021) and molecular analysis (Jumper et al., 2021). In order to explain how transformers emerged as the dominant neural architecture, it is important to understand the fundamental algorithmic primitives that transformers can solve in an efficient manner.
To that end, this paper addresses the following question.
Which architectural properties are necessary for transformers to solve graph algorithmic tasks?
In this work, we consider transformers that compute properties such as connectivity or the existence of a subgraph on some input graphs. We use this to develop a hierarchy of algorithmic tasks that proves sharp thresholds in the model width.
Graphs provide a compelling testbed for understanding transformer reasoning for serveral reasons. First, they serve as a natural “algorithmic playground,” encompassing a wide range of well-known problems that span computational classes. Many of these tasks have already been investigated as benchmarks for language models (Fatemi et al., 2023). Second, graph neural network (GNN) architectures (Gilmer et al., 2017; Kipf & Welling, 2017; Veličković et al., 2018; Hamilton et al., 2017) are a well-established class of models for graph-structured data, but they have well-known limitations on their expressive power (Xu et al., 2018; Loukas, 2019). This raises an important question about whether transformers are a remedy for those structural limitations of GNNs.
Finally, inference-time reasoning in large language models (LLMs), such as chain-of-thought (Wei et al., 2022) and graph-of-thought (Besta et al., 2024) prompting can often be framed as a graph search over some solution space. By investigating transformers through the lens of graph algorithms, we aim to deepen our understanding of their algorithmic capabilities and the role of design choices, such as model width and depth.
Questions about transformer algorithmic capabilities have attracted considerable attention, and several results that characterize what can and cannot be calculated using transformers of a certain size have emerged (e.g. Merrill & Sabharwal, 2023b; Liu et al., 2023). However, we are still far from a complete understanding of the interplay between architecture and algorithmic capabilities. In this work we focus on this problem in the context of graph algorithms, with emphases on the role of model depth and width and on a particular node-adjacency tokenization scheme.
Recent works (Sanford et al., 2024b, c) developed trade-offs between the abilities of transformers of different depths to solve fundamental graph problems.
They identified parallelism as a key aspect that distinguishes transformers, proving surprising benefits of employing a transformer rather than a GNN for graph-based tasks. For example, while GNNs require depth $O(n)$ to determine whether an input graph with $n$ vertices is connected, transformers can solve this task using only $O(\log n)$ layers of self-attention.
However, these results rely on restrictive assumptions about how graphs are represented as inputs to neural models. First, they tokenize an input graph as a sequence of discrete edge embeddings, rather than more standard node-based embeddings. Second, they focus on the scaling regime where the transformer’s embedding dimension (which we refer to interchangeably as its width) is restricted to be much smaller than the size of the input graph.
While a small width is certainly desirable, practical transformers have large widths with strict limitations on model depth.
For example, when learning over graphs, the width of the networks is often much larger than the number of vertices in the graph. Table 3 shows a list of commonly used graph datasets, including many molecular datasets, where the average graph size in the data is less than $40$. This is while commonly used network width is above $64$.
In this paper, we seek to understand which graph algorithms can be implemented in realistic transformer scaling regimes, such as when the width grows linearly in the number of tokens and the depth is constant, independent of the graph size. Concretely, we consider transformers that take as input graphs with adjacency embeddings for each node and place upper and lower bounds on the minimum width needed to solve several tasks. We focus on the constant-depth regime, where we do not allow the number of attention layers to scale as a function of the input size.
In Section 4, we introduce tasks—such as graph connectivity and fixed-length cycle detection—for which linear width is necessary and sufficient for dense graph inputs. Bounded-degree assumptions reduce the width threshold for the cycle detection task. Section 5 shows that more complex tasks—such as subgraph counting and Eulerian cycle verification—require super-linear width. We visualize the hierarchy over transformer widths induced by our collection of positive and negative results in Figure 1, which ranges from local node-level tasks that can be solved with constant width (such as computing the degree of each node), to arbitrary functions that necessitate a quadratic width scaling.
[…]
Figure 1: The width complexity hierarchy of graph tasks for transformers with node-adjacency tokenizations and constant depth. Each row visualizes the width regimes where the task is solvable (blue) or hard (red).
We validate our results empirically in Section 6. Specifically, we show that transformers with large embedding dimension and low depth achieve similar accuracy to deeper models, while having faster training and inference time, due to GPU parallelization.
2 Related Works
Expressive power of transformers
Transformer with arbitrary depth (Wei et al., 2021; Yun et al., 2020) or arbitrarily many chain-of-thought tokens (Malach, 2023) are known to be universal approximators. While these universality results are fundamental, they leave open questions about the ability of transformers to solve tasks efficiently, particularly in terms of parameter count and depth. To address this, various theoretical techniques have been employed to develop a more precise understanding of the scaling trade-offs involved.
For instance, Merrill & Sabharwal (2023a) show that constant-depth transformers with polynomial width can be simulated by $\mathsf{TC^{0}}$ Boolean circuits, implying that they cannot solve problems like graph connectivity. These results provide essential context for our paper; while we prove that transformers with a constant number of attention layers and bounded width can solve such tasks (e.g. in Theorem 4.1), this does not contradict their bounds, and together, they imply that a sufficiently large depth or width is necessary in the multi-layer perceptrons of our constructions. Similarly, Hao et al. (2022) identify formal language classes that can and cannot be recognized by hard-attention transformers. On the other hand, deeper transformers or those with sufficiently long chain-of-thought reasoning can simulate finite-state automata (Liu et al., 2023; Li et al., 2024).
A different perspective frames transformers in terms of communication complexity. Sanford et al. (2024c, b) draw an analogy between transformers and the Massively Parallel Computation (MPC) framework (Karloff et al., 2010), similar to prior work linking GNNs to the Congest model in distributed computing (Loukas, 2019). Sanford et al. (2024a) extend this analogy to define a transformer complexity hierarchy for graph tasks. While they focus on an edge-list tokenization and depth-based hierarchy, our work considers a width-based hierarchy with a node-adjacency tokenization, leading to different complexity insights.
Graph transformers.
Recent advancements in graph transformer architectures have significantly expanded the capabilities of Graph Neural Networks (GNNs) (Gilmer et al., 2017) by leveraging the power of attention mechanisms (Veličković et al., 2018; Brody et al., 2022) to facilitate dynamic node aggregation based on neighboring structures without relying on positional encodings. Building on this foundation, models like GraphBert (Zhang et al., 2020) incorporated graph-structured node embeddings to enhance representation learning. The evolution continued with the development of comprehensive graph transformer models that extend beyond traditional message-passing paradigms. For instance  Kreuzer et al. (2021) proposed graph transformers that generalize sequence-based transformers Subsequent models, including Spectral Attention Networks Kreuzer et al. (2021), introduced novel mechanisms for integrating spectral information and enhancing scalability for larger graphs. Additionally, GraphGPS (Rampášek et al., 2023) exemplifies the hybrid approach by combining attention mechanisms with message-passing techniques, thereby harnessing the strengths of both methods.
There is an abundance of possible embedding for graph transformers (Kreuzer et al., 2021; Dwivedi & Bresson, 2021; Kim et al., 2022; Rampášek et al., 2023; Zhang et al., 2024). Our work does not focus on comparing between such embeddings, but rather to study theoretically the effect of the embedding dimension. The embedding type itself is mainly a technical detail for proving our theoretical guarantees.
3 Problem setting and notations
3.1 Transformers
We consider the following setting of transformers: The input is a sequence of $N$ tokens $\mathbf{x}_{1},\dots,\mathbf{x}_{N}$ where $\mathbf{x}_{i}\in\mathbb{R}^{d_{\text{in}}}$. We denote by $X^{(0)}\in\mathbb{R}^{d_{\text{in}}\times N}$ the matrix where each row $i$ is equal to $\mathbf{x}_{i}$. Each layer of the transformer applies a self-attention mechanism on the inputs, and then an MLP. We denote by the input to layer $\ell$ as $X^{(\ell-1)}$. The self-attention at layer $\ell$ with $H$ heads is parameterized by matrices $K^{(\ell)}_{h},Q_{h}^{(\ell)}\in\mathbb{R}^{m_{\ell-1}\times m_{\ell-1}}$, $V^{(\ell)}_{h}\in\mathbb{R}^{m_{\ell-1}\times m_{\ell-1}}$. It is defined as:
$Z^{(\ell)}=\sum_{h=1}^{H}V^{(\ell)}_{h}X^{(\ell-1)}\text{sm}(X^{(\ell-1)^{\top}}K_{h}^{(\ell)^{\top}}Q^{(\ell)}_{h}X^{(\ell-1)})~{},$
where sm is row-wise softmax. The output is of dimension $Z^{(\ell)}\in\mathbb{R}^{m_{\ell-1}\times N}$. After that we use a residual connection, so the output of the self-attention layer is $\tilde{X}^{(\ell)}=Z^{(\ell)}+X^{(\ell-1)}$ Finally, we apply an MLP $\mathcal{N}^{(\ell)}:\mathbb{R}^{m_{\ell-1}}\rightarrow\mathbb{R}^{m_{\ell}}$ with ReLU activations on each token separately (i.e. each column of $\tilde{X}^{(\ell)}$). The output of the MLP is denoted as $X^{(\ell)}$, which is the input to the next layer. We denote by $\sigma:\mathbb{R}\rightarrow\mathbb{R}$ the ReLU activation.
We also often use positional encodings which are commonly used in standard transformer architectures. It is also common to add normalization layers, however it will not be necessary for our results, and these layers can be trivially added in all of our results without changing the output. The bit-precision of all our transformers will be $O(\log(n))$, where $n$ is the number of input tokens. This is a common assumption in many previous works (Sanford et al., 2024c, b; Merrill & Sabharwal, 2023b, 2024), and is relatively mild. We will denote $m=\max(m_{0},\dots,m_{L})$ the embedding dimension, and $d_{\text{in}}$ the input dimension.
Graph Inputs
Unlike GNNs, transformers are sequential architectures whose network topology does not encode the structure of an input graph. It is therefore a significant modeling decision to decide how a graph is best tokenized into a sequence of embeddings that can be proved to a transformer.
In this work, we focus primarily on the node-adjacency tokenization scheme, where each token corresponds to a node and encodes its incident edges.
Given a graph $G$ with $n$ nodes, let $A\in\mathbb{R}^{n\times n}$ be its adjacency matrix. The $i$th token input $\mathbf{x}_{i}\in\mathbb{R}^{n}$ to the transformer is defined as the $i$th row of $A$. Thus, $X\in\mathbb{R}^{n\times n}$.
There are numerous alternatives, but we restrict our focus to two others for experimental comparisons in Section 6. The edge-list tokenization of Sanford et al. (2024a) converts the graph into a sequence of discrete edge tokens. For a unique identifier assigned to each $v\in V$, this scheme encodes each edge $(v_{i},v_{j})\in E$ as a token $\mathbf{x}=(i,j)\in\mathbb{R}^{2}$.
A further alternative is the Laplacian eigenvector tokenization, which is introduced and discussed in Appendix A.
3.2 Graphs and Parallel Computation
Sanford et al. (2024b, a) introduced a representational equivalence between transformers with $L$ layers and $O(L)$ rounds of massively parallel computation (MPC). MPC (Karloff et al., 2010) is a theoretical formalization of the MapReduce (Dean & Ghemawat, 2008) distributed computing protocol. Under MPC, a very large number of machines with limited memory alternate between local computation and global communication rounds. The equivalence implies that transformers can act as MPC protocols by tokens as machines, local computation as MLP layers, and communication as self-attention. Critically, GNNs are not equivalent to MPC in the same sense, and are rather equivalent to the Local or Congest models, depending the sizes of each message between nodes is bounded (see Loukas, 2019). The main advantage of transformers over GNNs is the use of parallelism.
Graph algorithms are well-understood in the MPC framework. For instance, a wide range of graph tasks can be organized into depth-based equivalence classes that depend on the tasks’ computational complexity (Nanongkai & Scquizzato, 2022). While few unconditional negative results are known in the MPC framework, the one-cycle vs two-cycle task—which asks whether an input graph is either (1) a single cycle graph with $n$ nodes, or (2) a pair of cycle graphs, with $\frac{n}{2}$ nodes each—is widely expected to require either $\Omega(\log n)$ rounds or local memory of $\Omega(n)$. By combining these algorithms and conditional lower bounds with the transformer equivalence, we can obtain a numerous size complexity thresholds for transformers.
However, the results in Sanford et al. (2024b, a) have two limiting assumptions. The first is that the embedding dimension of the transformer
is strictly sublinear, namely $m=n^{1-\epsilon}$ for some constant $\epsilon\in(0,1)$ independent of $n$. The second limitation is that the graph is represented as an edge list, which requires $\Omega(n^{2})$ tokens for dense graphs. In this work we go beyond the strict sublinear regime, and focus on the node-adjacency tokenization strategy.
Remark 3.1.
When calculating the complexity of solving certain graph tasks, we usually only count the number of transformer layers, while the MLP can be arbitrarily large. This corresponds to the MPC model where usually only the distributed computing time is considered, while it is assumed that there is an unlimited local compute time for each machine. Intuitively, the number of self-attention layers is equivalent to the number of distributed computation rounds, while the MLP represent the local computation of each machine.
4 Beyond Sub-linear Embedding Dimension
In this section we show that having a transformer with linear embedding dimension can easily improve some known lower bounds based on distributed computing. We also show new lower bounds that are stricter for transformers with a linear embedding dimension, and a matching upper bound. All the proofs for the theorems in this section are in Appendix B.
4.1 Improving the sub-linear lower bounds
Previous works on the connection between the MPC model and transformers have shown conditional lower bounds for solving certain tasks. One such lower bound include the one-cycle versus two-cycles problem. Namely, given a graph with $n$ nodes, distinguish whether it is one cycle of length $n$ or two cycles of length $\frac{n}{2}$. Conjecture 13 from Sanford et al. (2024a) (see also (Ghaffari et al., 2019)) says that the MPC model with $n^{1-\epsilon}$ memory per machine for any $\epsilon\in(0,1)$ cannot distinguish between the two cases, unless the number of MPC rounds is $\Omega(\log(n))$. This also suggests that transformers with an embedding dimension of $O(n^{1-\epsilon})$ cannot solve this task. We will first show that having a transformer with linear embedding dimension can solve this task:
Theorem 4.1.
There exists a transformer with $2$ layers of self-attention, and embedding dimension $O(n)$ that solves the $1$ vs. $2$ cycle problem.
The proof intuition is that since the graph contains only $n$ edges, we can stack all of them into a single token. After doing so, we can use the MLP to solve it. This result shows that these kind of lower bounds for transformers are brittle in the sense that increasing the embedding dimension even slightly breaks them. With that said, this result is not surprising, since the graph in this task is very sparse.
The connectivity problem for general graphs cannot be solved in this manner.
The reason is that for graphs with $n$ nodes there are possibly $\Omega(n^{2})$ edges, and thus it is not possible to compress the entire graph into a single token with linear embedding dimension. However, the connectivity problem can still be solved by increasing the embedding dimension by polylog terms, as we explain below.
Ahn et al. (2012) uses linear sketching to solve the connectivity problem on general graphs with $O(n\log^{3}(n))$ total memory. The idea is to use linear sketching, which is a linear projection of the adjacency rows into vectors of dimension $O(\text{poly}\log(n))$. Although this is a lossy compression, it still allows to solve the connectivity problem. As a consequence, we can use a similar construction as in Theorem 4.1 where we first apply the sketching transformation to each token (i.e. row of the adjacency), then embed all the tokens, into a single token and use the MLP to solve the problem using the algorithm from Ahn et al. (2012) (Theorem 3.1).
Our next results will include problems that cannot be solved by simply compressing all the information of the graph into a single token, and would require a more intricate use of the self-attention layers.
4.2 The power of a linear embedding dimension
We have shown that transformers with a linear embedding dimension can improve certain lower bounds that apply for sub-linear dimensions. In this section we will show another lower bound on the task of detecting small cycles. This lower bound is stronger than the one for the one-cycle versus two-cycles in the sense that the depth-embedding dimension trade-off is even more strict.
Consider the task of detecting $2$-cycles in a directed graph. Namely, detecting whether there are nodes $u$ and $v$ with two edges, one directed from $u$ to $v$ and another directed from $v$ to $u$. We can prove the following lower bounds on this task:
Theorem 4.2.
Let $T$ be a transformer with embedding dimension $m$ depth $L$, bit-precision $p$ and $H$ attention heads in each layer. Also, assume that the input graphs to $T$ are embedded such that each token is equal to a row of the adjacency matrix. Then, if $T$ can detect $2$-cycles on directed graphs we must have that:
  1. 1.
    If $T$ has residual connections then $mpHL=\Omega(n)$.
  2. 2.
    If $T$ doesn’t have residual connections then $mpH=\Omega(n)$.
The proof uses a communication complexity argument, and specifically a reduction to set disjointness problem. For a formal definition see Appendix B. This lower bound is stronger than the lower bounds in Section 4.1 in the sense that they cannot be improved by having logarithmic depth.
For example, assume that the embedding dimension is $m=O(n^{1-\epsilon})$ for some $\epsilon\in(0,1)$, and that $p,H=O(\log(n))$ (which is often the case in practice). Then, transformers with residual connection require a depth of $\Omega(n^{\epsilon})$ to solve this task, which is beyond logarithmic, while transformer without residual connections cannot solve it. This lower bound is also unconditional, compared to the lower bouds from (Sanford et al., 2024a) which are conditional on the hardness of the one-cycle versus two-cycles problem (Conjecture 13 there). The caveat of this lower bound is that it relies on having the input graph specifically embedded as rows of the adjacency matrix.
We now turn to show an upper bound for this task. We will show an even more general claim than detecting $2$-cycles. Namely, that depth $L$ transformers with $\Omega(n)$ embedding dimension can calculate the $L$-th power of an adjacency matrix of graphs. In particular, the trace of $A^{L}$ provides the number of cycles of size $L$ in the graph (multiplied by an appropriate constant). Thus, a $2$-layer transformer with linear embedding dimension can already solve the directed $2$-cycle problem.
Theorem 4.3.
There exists an $O(L)$-layer transformer with embedding dimension $m=O(n)$ such that, for any graph embedded as rows of an adjacency matrix $A$, the output of the transformer in the $i$-th token is the $i$-th row of $A^{L}$.
The constructive proof of this theorem carefully selects key and query parameters to ensure that the output of the softmax matrix equals $A$. This enables an inductive argument that encodes $A^{\ell}$ in the value matrix of the $\ell$th layer, in order to compute $A^{\ell+1}$.
The above theorem shows that having a transformer with linear embedding dimension in the graph size is very strong in terms of expressivity. Namely, $A^{L}_{i,j}$ counts the number of walks of length $L$ form node $i$ to node $j$. This also allows to determine whether a graph is connected, by checking whether $A^{n}$ doesn’t contain any zero entries. Although, there are other algorithm (e.g. those presented in Section 4.1) that can solve the connectivity task more efficiently.
4.3 Sublinear embedding dimension for bounded-degree graphs
The results of the previous section establish that for worst-case graph instance, transformers with node-adjacency tokenizations require a linear embedding dimension to solve simple graph reasoning tasks, such as 2-cycle detection. This fact is perhaps unsurprising because each node-adjacency tokenization input is a length-$n$ boolean vector, which must be compressed in a lossy manner if the embedding dimension $m$ and bit precision $p$ satisfy $mp=o(n)$. It is natural to ask whether such results apply to sparse graphs as well, such as graphs with bounded degree.
Here, we show that requisite embedding dimension to detect 2-cycles scales linearly with the degree of the graph.
Theorem 4.4.
For any $n\in\mathbb{N}$ and $d\leq n$, there exists a single-layer transformer with embedding dimension $O(d\log n)$ that detects 2-cycles in any graph with node degree at most $d$. This embedding dimensional is optimal up to logarithmic factors.
The proof reduces the dimensionality of the input adjacency matrix by incorporating vector embeddings from (Sanford et al., 2024c) into the key, query, and value matrices to produce a “sparse attention unit,” whose activations are large when a respective cycle exists.
5 Super-linear embedding dimension
In the previous section we have shown the power of having a transformer with an embedding dimension that is linear in the graph size. In this section we study transformers with a super-linear dimension.
Permitting the embedding dimension to $\Omega(n^{2})$ trivializes the question of expressive power in our setting.
In this regime, the entire graph can be encoded into a single token, which offsets any algorithmic demands to the computationally-unbounded MLP.
As a result, we restrict this section’s focus to embedding dimensions that are super-linear, but sub-quadratic, in the number of nodes.
5.1 Detecting and counting subgraphs
A very basic graph task is to detect or count a given subgraph. This task is important in several fields including biology, organic chemistry, graph kernels and more (See (Jin et al., 2020; Jiang et al., 2010; Pope et al., 2018; Shervashidze et al., 2009) and the discussion in Chen et al. (2020)). This task was also studied in the context of GNNs in Chen et al. (2020), where it was shown that $k$-IGN can detect subgraphs of size $k$. The caveat of this result is that $k$-IGN require $k$-tensor to construct, which is practically infeasible even for small $k$.
Here we study the ability of transformer to detect and count subgraphs. Our main result is the following:
Theorem 5.1.
Let $k,n\in\mathbb{N}$, and let $G^{\prime}$ be a graph with $k$ nodes. There exists a transformer with $O(1)$ self-attention layers and embedding dimension $O\left(n^{2-1/k}\right)$ such that for any graph $G$ of size $n$,
counts the number of occurrences of $G^{\prime}$ as a subgraph of $G$.
The crux of the proof is to implement the seminal ”tri-tri-again“ algorithm (Dolev et al., 2012) using transformers. The general idea is that given a graph with $n$ nodes, we split the nodes into $n^{1/k}$ disjoint sets, each containing $n^{1-1/k}$ nodes, where $k$ is the size of the subgraph that should be detected or counted. For each possible combination of $k$ such sets we use an MLP to count the given subgraph in it. There are $\binom{n^{1/k}}{k}\leq n^{k\cdot 1/k}=n$ such combinations of sets, each one contains at most $n^{2-2/k}$ edges. Thus, each token with a large enough embedding dimension can simulate one combination of subsets, cumulating all the relevant edges. Note that subgraph detection is a sub-task of counting, whether the number of occurrences is larger than 00.
Theorem 5.1 can be compared to Theorem 23 in Sanford et al. (2024a) that provides a construction for counting $k$-cliques using transformers with sub-linear memory and additional blank tokens. There, it was shown that it is possible to count $k$-cliques with a transformer of depth $O(\log\log(n))$, however the number of blank tokens is $O(n^{k-1})$ in the worse case. By blank tokens we mean additional empty tokens that are appended to the input and used for scratch space as defined in (Sanford et al., 2024a). In our result, we require a depth of $O(1)$, and the total number of tokens is $n$, while the embedding dimension is super-linear (but sub-quadratic). Thus, our solution has better memory usage, since the increase in width is only polynomial, but the number of tokens is $n$ instead of $O(n^{k})$.
Remark 5.2.
We note that the overall computation time of our depth $O(1)$ model above is still exponential in $k$ due to the size of the MLP. This is also the case in Theorem 23 from Sanford et al. (2024a). In fact, it is a common conjecture in the literature that no (possibly randomized) algorithm can detect $k$-cliques in time less than $O(n^{2k/3-\epsilon})$ for any $\epsilon>0$ (see Hypothesis 6 in Williams (2018)).
5.2 Limitations of sub-quadratic memory
We now turn to the boundary of the possible embedding dimension for which these kinds of questions are interesting. A natural question is whether there is a width that suffices for solving all graph problems. Intuitively, this width should be n2 as this is what is required for storing the entire graph. Here we show that this is indeed an upper bound on needed depth for an arbitrary problem, and that this bound is tight for a particular problem: the Eulerian cycle verification problem.
Given a graph and a list of “path fragments,” where each fragment consists of a pair of subsequent edges in a path, the goal of the Eulerian cycle verification problem is to determine whether the properly ordered fragments comprise an Eulerian cycle. Recall that a Eulerian cycle is a cycle over the entire graph that uses each edge exactly once.
We show that this problem on multigraphs with self loops cannot be solved (conditioned on the $1$ vs. $2$ cycle conjecture) with a transformers unless its embedding dimensions is quadratic in $n$ or its depth is $\log(n)$.
Theorem 5.3.
Under Conjecture 2.4 from Sanford et al. (2024b), the Eulerian cycle verification problem on multigraphs with self loops cannot be solved by transformers with adjacency matrix inputs if $m=O\left(n^{2-\epsilon}\right)$ for any constant $\epsilon>0$, unless $L=\Omega(\log(n))$.
Figure 2: Training and inference times (a), training loss curves (b), and accuracy curves (c) for the connectivity task over graphs with $100$ nodes, across transformers with approximately 100k parameters, varying in width and depth.
While the loss and accuracies remain consistent, shallow and wide transformers demonstrate significantly faster training and inference times.
6 Experiments
In this section, we present experiments to support our claims and theoretical results.  [^1] Code is provided in the Supplementary Material. In our experiments, we used a standard transformer architecture using Pytorch’s transformer encoder layers (Paszke et al., 2019). Specifically, each layer is composed of Multi-Head Self-Attention, Feedforward Neural Network, Layer Normalization and Residual Connections. More experimental details are provided in Appendix E.
For all experiments in Section 6.1 and Section 6.2, we used the adjacency rows tokenization as described in Section 5.1. Details of the implementation are described in Appendix D. We considered the tasks of connectivity, triangles count, and 4-cycle count. For the counting tasks, we used the substructure counting dataset from Chen et al. (2020), where each graph was labeled with the number of pre-defined substructures it contains, as a graph regression task. For the connectivity tasks, we generated synthetic graphs, and the label indicates whether the graph is connected or not. All the datasets information is described in detail in  Appendix E.
6.1 Empirical Trade-Offs Between Width and Depth
In this subsection, we examine the empirical trade-offs between depth and width in transformers. We show that when using transformers that are shallow and wide, the training and inference times are significantly lower than when using deep and narrow transformers. This is while test error and training convergence rates are empirically similar.
We trained a transformer with a fixed amount of 100k parameters split between varying depth and width. We examine how the running time, loss, and generalization depend on the width and depth. We examine the following pairs of (depth, width): $(1,125),(2,89),(4,63),(8,45),(10,40)$. We train each model for $100$ epochs and examine the following: the total training time, the total inference time, the training loss and test performance (accuracy for classification and MAE for regression). We repeat this experiment with graph sizes 50 and 100. We report the averages over $3$ runs with random seeds. The hyper-parameters we tuned are provided in Appendix E.
The results for the connectivity task over $100$ nodes are presented in Figure 2. Additional results for the counting tasks and different graph sizes, present the same trends, and are provided in Appendix D due to space limitations. As shown in Figure 2 and Figure 2, the training loss and accuracy remain consistent across all depth and width configurations. However, Figure 2 reveals that shallow and wide transformers significantly reduce the total training and inference time compared to their deeper and narrower counterparts. This may be due to the ability of GPUs to parallelize the computations across the width of the same layer, but not across different layers.
6.2 Critical Width
In this subsection, we show that the transformer width required to fit the substructure counting tasks increases roughly linearly with the number of nodes, as argued in  Section 5.1.
The experiments considered graphs with increasing numbers of nodes, ranging from 50 to 400 in increments of 50, and transformer widths varying from 100 to 800 in increments of 100. For each combination of graph size and transformer width, we determined the critical width at which the model failed to fit the data. The critical width is defined as the width where the training loss plateaued at more than 0.05. To determine the critical width, we conducted a grid search over each combination of graph size and model width, and selected the model that fitted the data best.
The hyper-parameters we considered are provided in Appendix E.
To isolate the effect of width, we used a transformer model with one layer. We used two attention heads to ensure there exists a width for which the model can fit the data for all the evaluated graph sizes. The results for the 4-Cycle count task are presented in Figure 3. Due to space limitations, the results for the triangle count task are deferred to Appendix D. This figure shows that the critical width increases roughly linearly with the graph size.
Figure 3: Critical width evaluation for the 4-Cycle Count task, using a transformer with 1 layer. The points indicate the critical width at which the model fails to fit the data.
6.3 Adjacency Rows Graph Tokenization
In this paper, we focus on a graph tokenization where each row of the graph’s adjacency matrix is treated as a token for the model. This tokenization offers significant efficiency advantages for dense graphs, as the edge-list representation requires $O(n^{2})$ tokens, whereas the adjacency-row representation reduces this to $O(n)$. To validate the effectiveness of this tokenization approach in practice, we evaluate it on real graph datasets.
We compared the adjacency-row representation to the edge-list representation by training a transformer model on three Open Graph Benchmark (OGB) (Hu et al., 2020) datasets: ogbg-molhiv, ogbg-molbbbp, and ogbg-molbace.
In ogbg-molhiv, the task is to predict whether a molecule inhibits HIV replication, a binary classification task based on molecular graphs with atom-level features and bond-level edge features.
ogbg-bbbp involves predicting blood-brain barrier permeability, a crucial property for drug development.
ogbg-bace focuses on predicting the ability of a molecule to bind to the BACE1 enzyme, associated with Alzheimer’s disease.
We also evaluated a tokenization using the eigenvectors and eigenvalues of the graph Laplacian, as commonly used in the literature (Dwivedi & Bresson, 2021; Kreuzer et al., 2021). More experimental details, including the dataset statistics can be found in Appendix E.
The results of our evaluation are summarized in Table 1. In all three tasks, the adj-rows representation achieved better ROC-AUC scores than the edge-list representation.
In two out of the three, it also improved upon the commonly used Laplacian eigenvectors representations.
The results suggest that the adjacency representation we use in this paper is empirically effective, and should be considered alongside the commonly used Laplacian eigenvector representation.
Table 1: This table compares the ROC-AUC performance metrics for different graph representations: Edge List, Adjacency Rows, and Laplacian Eigenvectors (LE).
The results are reported as the average ROC-AUC over three random seeds, along with the corresponding standard deviations.
Dataset
Tokenization molhiv molbbbp molbeca
EdgeList 54.01±1.38 64.73±1.66 66.06±3.89
AdjRows 61.87±1.10 67.63±2.57 68.64±2.34
LE 68.11±1.52 55.31±4.79 63.61±2.31
7 Discussion and Future Work
This paper uses a collection of graph algorithmic tasks—including subgraph detection, connectivity, and Eulerian cycle verification—to demonstrate the powers of width bounded-depth transformers that take as input node adjacency encodings. These results include sharp theoretical thresholds that demonstrate the trade-offs between constant, linear, quadratic, and intermediate width regimes. Our empirical results validate the efficiency and accuracy of our choice of scaling regime and embedding strategy.
There are numerous possible extensions of this work.
One future direction is to study different graph tokenization schemes, beyond the node-adjacency encoding of this work and edge-list encoding of (Sanford et al., 2024a).
A particularly notable alternative is the smallest eigenvectors of the graph Laplacian, presented as a vector of components for each node.
This spectral embedding is a standard embedding scheme for GNNs, and the techniques and tasks developed in this paper would likely be relevant to proving similar bounds.
We provide a preliminary exploration of the trade-offs between the intrinsically local characteristics of adjacency-based tokenization schemes and more global spectral approaches in Appendix A.
Another future direction is to study the optimization and generalization capabilities of transformers to solve graph problems, beyond the expressiveness limitations presented in this work.
Acknowledgments
This work was supported in part by the Tel Aviv University Center for AI and Data Science (TAD), the Israeli Science Foundation grants 1186/18 and 1437/22. OF was partially supported by the Israeli Science Foundation (grant No. 1042/22 and 800/22). We thank Joan Bruna for the helpful discussions while this work was being completed.
Appendix A Alternative tokenization approaches
While the primary aim of this paper is to study the properties of the node-adjacency tokenization scheme in terms of its width and depth trade-offs, we also establish clear trade-offs between this scheme and other encoding schemes. The Laplcian-eigenvector tokenization passes as input to the transformer each node’s components of the most significant eigenvectors.
Concretely, let $A\in\mathbb{R}^{n\times n}$ be the adjacency matrix of a graph and $D$ the diagonal degree metrix. The Laplacian matrix is defined as $\mathcal{L}=D-A$. Denote the eigenvectors of $\mathcal{L}$ as $\mathbf{v}_{1},\dots,\mathbf{v}_{n}$ with respective eigenvalues $0=\lambda_{1}\leq\dots\leq\lambda_{n}$. For some embedding dimension $m$, we let the $m$-dimensional Laplacian-eigenvector tokenization be $\mathbf{y}_{1},\dots\mathbf{y}_{n}$, where $\mathbf{y}_{i}=(\mathbf{v}_{1,i},\dots,\mathbf{v}_{m,i})$; we encode the eigenvalues as well as $\mathbf{y}_{0}=(\lambda_{1},\dots,\lambda_{m})$. We contrast this with node-adjacency encodings of embedding dimension $m$, whose $i$th input is $\mathbf{x}_{i}=\phi(A_{i})$.
We note several illustrative toy tasks that demonstrate trade-offs between the two graph tokenization schemes.
Node-adjacency advantage at local tasks
The node-adjacency tokenization is amenable for analyzing local structures around each node. Most simply, the degree of each node can be computed in a sequence-wise manner with node-adjacency tokenization with embedding dimension $m=1$ by simply computing the inner products $\langle\boldsymbol{1}_{n},\mathbf{x}_{i}\rangle$. Constructions like Theorems 4.3 and 4.4 further demonstrate the abilities of adjacency encodings to aggregate local structures.
In contrast, choosing the smallest eigenvectors in the alternative encoding makes it impossible to even compute each node degree without having embedding dimension $m$ growing linearly in the node count $n$222 Consider the task of computing the degree of a particular node of a graph consisting of $\frac{n}{3}$ disconnected linear subgraphs, each with three nodes connected by two edges. The zero eigenvalue thus has multiplicity $\frac{n}{3}$, and hence the eigenvectors $\mathbf{v}_{1},\dots,\mathbf{v}_{n/3}$ exist solely as indicators of connected components. Therefore, if nodes $i,j,k$ comprise a cluster and $m\leq\frac{n}{3}$, then their embeddings $\mathbf{y}_{i},\mathbf{y}_{j},\mathbf{y}_{k}$ are identical. .
Laplacian-eigenvector advantage at global tasks
In contrast, the most significant graph Laplacian provide high-level information about the global structure of the graph. Most notably, the tokenization trivializes the connectivity task because a graph is disconnected if and only if its second-smallest eigenvalue is zero; transformers with the node-adjacency tokenization require either depth $\Omega(\log n)$ or width $\Omega(n)$ to solve the same problem.
Other properties of structured graphs reveal themselves with low-dimensional Laplacian-eigenvector tokenizations. For instance, the relative position of a node in a lattice or ring graph are encoded in the most significant eigenvectors. Graph clustering algorithms could be inferred by transformers that take spectral encodings as input and simulate algorithms like $k$-means. The hardness of graph connectivity with the adjacency encoding translates to hardness results for efficiently simulating clustering algorithms.
Quadratic embedding equivalence
Critically, the above trade-offs occur in small embedding dimensions. In the regime where $m=\Omega(n^{2})$ and MLPs are universal approximators, both tokenization schemes are universal. The entire graph can be encoded in a single token, which can then convert between $A$ and the spectrum of $\mathcal{L}$.
Appendix B Proofs from Section 4
B.1 Proof of Theorem 4.1
See 4.1
Proof.
The proof idea is to embed all the information about the graph into a single token, and then offload the main bulk of the solution to the MLP. For that, the first layer will transform the input of each node from adjacency rows to only indicate its two neighbors. The second layer will embed all the information over the entire graph into a single token.
We now define the construction of the transformer. The input to the transformer are adjacency rows, where we concatenate positional encodings that include the row number. Namely, the $i$-th input token is equal to $\begin{pmatrix}\mathbf{x}_{i}\\ i\end{pmatrix}$, where $\mathbf{x}_{i}$ is the $i$-th row of the adjacency matrix of the graph. The first layer of self-attention will not effect the inputs. This can be thought of as choosing $V=\boldsymbol{0}$ (while $K$ and $Q$ are arbitrary), and using the residual connection so that the tokens remain the same as the input. We now use Lemma B.1 to construct a $3$-layer MLP that changes the embedding of each token such that it includes for each node its neighbors. The MLP will does not change the positional encoding, this can be done since ReLU network can simulate the identity function by $z\mapsto\sigma(z)-\sigma(-z)$. We add another layer to the MLP that maps $\mathbb{R}^{3}\ni\mathbf{v}_{i}\mapsto\mathbf{u}_{i}\in\mathbb{R}^{3n}$, where $(\mathbf{u}_{i})_{3(i-1)+1:3i}=\mathbf{v}_{i}$ and all the other entries of $\mathbf{u}_{i}$ are equal to 00.
The second layer of self-attention will have the following matrices: $K=Q=\boldsymbol{0}_{3n\times 3n},V=n\cdot I_{3n}$. Since we used the zero attentions, all the tokens attend in the same way and using the exact same weight to all other tokens. The Softmax will normalize the output by the number of tokens, namely by $n$. Hence, after applying the $V$ matrix, all the output tokens of the second layer of self-attention will be equal to the sum of all the tokens that were inputted to the second layer.
In total, we get that the output of the second layer of attention is a vector with $3n$ coordinates, where each $3$ coordinates of the form $\begin{pmatrix}i\\ j\\ k\end{pmatrix}$ represent the two edge $(i,k),(j,k)$. Thus, the entire information of the graph is embedded in this vector.
Finally, we use the MLP to determine whether the input graph, whose embedded as a list of edges, is connected. This can be done by an MLP since it has the universal approximation property (Cybenko, 1989; Leshno et al., 1993). Although we don’t specify the exact size of this MLP, it can be bounded since there are efficient deterministic algorithms for determining connectivity. These algorithms can be simulated using ReLU networks.
Note that the output of the connectivity problem is either 00 or $1$, thus it is enough to approximate a solution of this task up to a constant error (say, of $\frac{1}{4}$), and then use another layer to threshold over the answer.
Lemma B.1.
There exists a $3$-layer MLP $\mathcal{N}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{2}$ such that for every vector $\mathbf{v}$ where there are $i,j\in[n]$ with $(\mathbf{v})_{i}=(\mathbf{v})_{j}=1$ and $(\mathbf{v})_{k}=0$ for every other entry we have that either $\mathcal{N}(\mathbf{v})=\begin{pmatrix}i\\ j\end{pmatrix}$ or $\mathcal{N}(\mathbf{v})=\begin{pmatrix}j\\ i\end{pmatrix}$.
Proof.
The first layer of the MLP will implement the following function:
$\mathbf{v}\mapsto\begin{pmatrix}\sum_{i=1}^{n}i\cdot\mathds{1}((\mathbf{v})_{i}=1)\\ \sum_{i=1}^{n}i^{2}\cdot\mathds{1}((\mathbf{v})_{i}=1)\end{pmatrix}~{}.$
This is a linear combination of indicators, where each indicator can be implemented by the function $f(z)=\sigma(x)-\sigma(x-1)$. Note that the input to the MLP is either 00 or $1$ in each coordinate, thus the output of this function will be $\begin{pmatrix}i+j\\ i^{2}+j^{2}\end{pmatrix}$. We have that $i+j$ and $i^{2}+j^{2}$ determine the values of $i$ and $j$. This means that if $i_{1}+j_{1}=i_{2}+j_{2}$ and $i_{1}^{2}+j_{1}^{2}=i_{2}^{2}+j_{2}^{2}$ then $i_{1}=i_{2}$ or $i_{1}=j_{2}$ and similarly for $i_{2}$. Since there are $O(n^{2})$ different possible values, we can construct a network with $2$-layers and $O(n^{2})$ width that outputs $\begin{pmatrix}i\\ j\end{pmatrix}$ up to changing the order of $i$ and $j$.
B.2 Proof of Theorem 4.2
See 4.2
Proof.
Our proof relies on a communication complexity lower bound for the set disjointness problem, and is similar to the arguments from (Sanford et al., 2024c; Yehudai et al., 2024). The lower bound for communication complexity is the following: Alice and Bob are given inputs $a,b\in\{0,1\}^{s}$ respectively, and their goal is to find $\max a_{i}b_{i}$ by sending single bit-messages to each other in a sequence of communication rounds. The lower bound says that any deterministic protocol for solving such a task requires at least $s$ rounds of communication.
We set $s=n^{2}$, and design a graph $G=(V,E)$ that has a directed $2$-cycle iff $\max a_{i}b_{i}=1$. The graph has $|V|=2n$, we partition the vertices into $2$ disjoint sets $V_{1},V_{2}$, and number the vertices of each set between $1$ and $n$. The inputs $a$ and $b$ encode the adjacency matrices between vertices in $V_{1}$ and $V_{2}$, and between vertices in $V_{2}$ and $V_{1}$ respectively. Now, there exists a directed $2$-cycle iff there is some $i\in[s]$ for which both $a_{i}=1$ and $b_{i}=1$ meaning that $\max a_{i}b_{i}=1$.
Assume there exists a transformer of depth $L$ with $H$ heads, embedding dimension $m$ and bit precision $p$ that successfully detects $2$-cycles in a directed graph. Denote the weights of head $i$ in layer $\ell$ by $Q_{i}^{\ell},K_{i}^{\ell},V_{i}^{\ell}\in\mathbb{R}^{m\times m}$ for each $i\in[H]$, and assume w.l.o.g. that they are of full rank, otherwise our lower bound would include the rank of these matrices instead of the embedding dimension (which can only strengthen the lower bound). We design a communication protocol for Alice and Bob to solve the set disjointness problem. The communication protocol will depend on whether $T$ has residual connections or not. We begin with the case that it does have them, the protocol works as follows:
  1. 1.
    Given input sequences $a,b\in\{0,1\}^{s}$ to Alice and Bob respectively, they calculate the input tokens $x_{1}^{0},\dots,x_{n}^{0}$ and $x_{n+1}^{0},\dots,x_{2n}^{0}$, respectively. Note that the adjacency matrix have a block shape, thus both Alice and Bob can calculate the rows of the adjacency matrix corresponding the the edges which are known to them.
  2. 2.
    Bob calculates $K_{j}^{1}x_{i}^{0},Q_{j}^{1}x_{i}^{0},V_{j}^{1}x_{i}^{0}$ for every head $j\in[H]$ and transmits them to Alice. The number of transmitted bits is $O(nmHp)$
  3. 3.
    Alice can now calculate the output of the $r$-th token after the first layer. Namely, for every head $j\in[H]$, she calculates:
    $s_{j}^{r}=\sum_{i=1}^{2n}\exp(x_{i}^{0^{\top}}K_{j}^{1^{\top}}Q_{j}^{1}x_{r}^{0})$
    $t_{j}^{r}=\sum_{i=1}^{2n}\exp(x_{i}^{0^{\top}}K_{j}^{1^{\top}}Q_{j}^{1}x_{r}^{0})V_{j}^{1}x_{i}^{0}~{}.$
    The output of the $j$-th head on the $r$-th token is equal to $\frac{t_{j}^{r}}{s_{j}^{r}}$. For the first $n$ tokens, Alice use the residual connection which adds the tokens that are known only to her. She now passes the tokens through the MLP to calculate $x_{1}^{1},\dots,x_{n}^{1}$, namely the output of the tokens known to her after the first layer.
  4. 4.
    Similarly to the previous $2$ steps, Bob calculates the tokens $x_{n+1}^{1},\dots,x_{2n}^{1}$ which are known only to him.
  5. 5.
    For any additional layer, the same calculations are done so that Alice calculates $x_{1}^{\ell},\dots,x_{n}^{\ell}$ and Bob calculates $x_{n+1}^{\ell},\dots,x_{2n}^{\ell}$.
In case there are no residual connections, after the third step above Alice have the information about all the tokens. Hence, there is no need for more communication rounds, and Alice can finish the rest of the calculations of the transformers using the output tokens of the first layer.
By the equivalence between the set disjointness and the directed $2$-cycle that was described above, Alice returns $1$ iff the inputs $\max_{i}a_{i}b_{i}=1$, and 00 otherwise. The total number of bits transmitted in this protocol in the case there are residual connection is $O(nmpHL)$, since there are $O(nmpH)$ bits transferred in each layer. The lower bound is determined by the size of the input which is $s=n^{2}$, hence $mpHL=\Omega(n)$. In the case there are no residual connections there is no dependence on $L$, hence the lower bound becomes $mpH=\Omega(n)$. ∎
B.2.1 An extension to bounded degree graphs
In order to prove the optimality result in Theorem 4.4 for the task of determining the existence of a 2-cycle in bounded-degree graphs, we state the following theorem.
Theorem B.2.
Let $T$ be a transformer with embedding dimension $m$ depth $L$, bit-precision $p$ and $H$ attention heads in each layer. If $T$ can detect $2$-cycles on $d$-degree directed graphs, then:
  1. 1.
    If $T$ has residual connections then $mpHL=\Omega(d)$.
  2. 2.
    If $T$ doesn’t have residual connections then $mpH=\Omega(d)$.
Proof.
The proof is nearly identical to that of Theorem 4.2, except that we alter the reduction to ensure that that the graph possessed by Alice and Bob is of degree $d$.
As before, we design a graph $G=(V,E)$ with vertices partitioned into two sets $V_{1},V_{2}$ satisfying $|V_{1}|=|V_{2}|=n$. Let $\bar{E}_{d}$ denote the edges of a bipartite graph between $V_{1}$ and $V_{2}$ such that (1) every node has $d$ incident outgoing edges; and (2) $(i,j)\in\bar{E}_{d}$ if and only if $(j,i)\in\bar{E}_{d}$.
Consider some instance of set disjointness with $a,b\in\{0,1\}^{s}$ for $s=nd$. We index the $2s$ edges in $\bar{E}_{d}$ as $e^{a}_{1}=(v^{1}_{1},v^{2}_{1}),\dots,e^{b}_{s}=(v^{1}_{s},v^{2}_{s})$ and $e^{b}_{1}=(v^{2}_{1},v^{1}_{1}),\dots,e^{b}_{s}=(v^{2}_{s},v^{2}_{s})$. Then, we embed the instance by letting $e^{a}_{i}\in E$ if $a_{i}=1$ and $e^{b}_{i}\in E$ if $b_{i}=1$. As before, there exists a directed 2-cycle in $G$ if and only if $\max_{i}a_{i}b_{i}=1$.
The analysis of the transformer remains unchanged. An $L$-layer transformer with embedding dimension $m$, heads $H$, and bit precision $p$ transmits $O(nmpHL)$ bits between Alice and Bob. The hardness of set disjointness requires that at least $s=nd$ bits be transmitted, which means that it must be the case that $mphL=\Omega(d)$. ∎
B.3 Proof of Theorem 4.3
We will need the following lemma for the proof.
Lemma B.3.
Let $a_{1},\dots,a_{k},b_{1},\dots,b_{k}\in\mathbb{R}$ where the $a_{i}$’s are distinct. There exists a $2$-layer fully-connected neural network $N:\mathbb{R}\rightarrow\mathbb{R}$ with width $O(k)$ such that $N(a_{i})=b_{i}$.
Proof.
Let $\delta=\min_{i}\in[k]|a_{i}-a_{j}|$, by the assumption $\delta>0$. Let:
$f_{i}(x)=\frac{1}{\delta}\left(\sigma(x-(a_{i}-2\delta))-\sigma(x-(a_{i}-\delta)+\sigma(a_{i}+2\delta-x)-\sigma(a_{i}+\delta-x)\right)~{}.$
It is clear that $f_{i}(a_{i})=1$ and $f_{i}(a_{j})=0$ for any $j\neq i$. Thus, we define the network: $N(x)=\sum_{i=1}^{k}b_{i}f_{i}(x)$. ∎
We are now ready to prove the theorem.
See 4.3
Proof.
We will first show the construction for the case of $L=1$, and then show inductively for general $L$. We define the input for the transformer as $X=\begin{pmatrix}A\\ I\\ \boldsymbol{0}_{d\times d}\\ \boldsymbol{0}_{d\times d}\end{pmatrix}\in\mathbb{R}^{3d\times d}$, namely, there are $d$ tokens, each token contains a column of the adjacency matrix concatenated with a positional embedding. The self-Attention layer contains one head with the following matrices:
$K=c\cdot\begin{pmatrix}I&&\\ &\boldsymbol{0}_{d\times d}&\\ &&\boldsymbol{0}_{d\times d}\end{pmatrix},~{}~{}Q=\begin{pmatrix}\boldsymbol{0}_{d\times d}&I&\\ \boldsymbol{0}_{d\times d}&\boldsymbol{0}_{d\times d}&\\ &&\boldsymbol{0}_{d\times d}\end{pmatrix}~{},V=\begin{pmatrix}I&&\\ &I&\\ &&\boldsymbol{0}_{d\times d}\end{pmatrix}~{}.$
where $c>0$ is some sufficiently large constant the determines the temperature of the softmax. We first have that $X^{\top}K^{\top}QX=A$. Since all the values of $A$ are either 00 or $1$, for a sufficiently large $c>0$, the softmax behave similarly to the hardmax and we get that:
$VX\text{sm}(A)=\begin{pmatrix}A^{2}\cdot\text{deg}(A)^{-1}\\ A\cdot\text{deg}(A)^{-1}\\ \boldsymbol{0}_{d\times d}\end{pmatrix}~{},$
where $\text{deg}(A)$ is a diagonal matrix, where its $i$-th diagonal entry is equal to the degree of node $i$. Finally, we apply an MLP $\mathcal{N}:\mathbb{R}^{3d\times d}\rightarrow\mathbb{R}^{3d\times d}$ that operates on each token separately. We define the MLP such that:
$\mathcal{N}\left(\begin{pmatrix}A^{2}\cdot\text{deg}(A)^{-1}\\ A\cdot\text{deg}(A)^{-1}\\ \boldsymbol{0}_{d\times d}\end{pmatrix}\right)=\begin{pmatrix}\boldsymbol{0}_{d\times d}\\ \boldsymbol{0}_{d\times d}\\ A^{2}\end{pmatrix}~{}.$
Constructing such an MLP can be done by calculating the degree of each token from $A\cdot\text{deg}(A)^{-1}$ and multiplying the first $d$ coordinates of each token by this degree. This can be done since the entries of this matrix is either 00 or the inverse of the degree of node $i$, thus it requires only inverting an integer between $1$ and $n$. By Lemma B.3 this can be done by a $2$-layer MLP with width $n$. This finishes the construction for calculating $A^{2}$.
For general $L>2$ we use the residual connection from the inputs. That is, the input to the $L$-th layer of the transformer is equal to $\begin{pmatrix}A\\ I\\ A^{L}\end{pmatrix}$. We use a similar construction as the above, except that we use the matrix $V=\begin{pmatrix}\boldsymbol{0}_{d\times d}&&\\ &I&\\ &&I\end{pmatrix}$. This way, the output of the self-attention layer is $\begin{pmatrix}\boldsymbol{0}_{d\times d}\\ A\cdot\text{deg}(A)^{-1}\\ A^{L+1}\cdot\text{deg}(A)^{-1}\end{pmatrix}$, and we employ a similar MLP as before to eliminate the $\text{deg}(A)^{-1}$ term. ∎
B.4 Proof of Theorem 4.4
See 4.4
The optimality result is proved using the same methodology as Theorem 4.2. It is stated and proved formally as Theorem B.2.
The proof of the construction adapts an argument from Theorem 2 of (Sanford et al., 2024c), which shows that a sparse averaging task can be solved with bounded-width transformers. We make use of the following fact, which is a consequence of the Restricted Isometry Property analysis of (Candes & Tao, 2005; Mendelson et al., 2005).
Lemma B.4.
For any $d\leq n\in\mathbb{N}$ and $p=\Omega(d\log n)$, there exist vectors $\mathbf{y}_{1},\dots,\mathbf{y}_{n}\in\mathbb{R}^{p}$ such that for any $\mathbf{x}\in\{0,1\}^{n}$ with $\sum_{i}\mathbf{x}_{i}\leq d$, there exists $\phi(\mathbf{x})\in\mathbb{R}^{p}$ such that
$\langle\phi(\mathbf{x}),\mathbf{y}_{i}\rangle$ $=1,\quad\text{if $\mathbf{x}_{i}=1$},$
$\langle\phi(\mathbf{x}),\mathbf{y}_{i}\rangle$ $\leq\frac{1}{2},\quad\text{if $\mathbf{x}_{i}=0$.}$
We use this fact to prove Theorem 4.4.
Proof.
Concretely, we prove that some transformer $T$ exists that takes as input
$X=\begin{pmatrix}\mathbf{x}_{1}&\dots&\mathbf{x}_{n}\\ 1&\dots&n\end{pmatrix}\in\mathbb{R}^{(n+1)\times n}$
and returns $T(X)\in\{0,1\}^{n}$, where $T(X_{i})=1$ if and only if the $i$th node in the graph whose adjacency matrix is $A$ belongs to a directed 2-cycle. We assume that no self-edges exist.
We first configure the input MLP to incorporate the above vectors for node identifiers and adjacency rows and produce the tokens $\tilde{X}=(\tilde{\mathbf{x}}_{1},\dots\tilde{\mathbf{x}}_{n})\in\mathbb{R}^{m\times n}$ for $m=2p+2$:
$\tilde{\mathbf{x}}_{i}=\begin{pmatrix}\phi(\mathbf{x}_{i})\\ \mathbf{y}_{i}\\ 1\\ 0\end{pmatrix}.$
We also introduce a constant-valued “dummy node” $\tilde{\mathbf{x}}_{n+1}$, which has no edges and does not appear in the output:
$\tilde{\mathbf{x}}_{n+1}=\begin{pmatrix}\boldsymbol{0}_{p}\\ \boldsymbol{0}_{p}\\ 0\\ 1\end{pmatrix}.$
We define linear transforms $Q,K,V\in\mathbb{R}^{d\times n}$ that satisfy the following, for any $i\in[n]$ and some sufficiently large temperature constant $c$:
$Q\tilde{\mathbf{x}}_{i}=c\begin{pmatrix}\phi(\mathbf{x}_{i})\\ \mathbf{y}_{i}\\ \frac{7}{4}\\ 0\end{pmatrix},\ K\tilde{\mathbf{x}}_{i}=\begin{pmatrix}\mathbf{y}_{i}\\ \phi(\mathbf{x}_{i})\\ 0\\ 0\end{pmatrix},\ K\tilde{\mathbf{x}}_{n+1}=\begin{pmatrix}\boldsymbol{0}_{p}\\ \boldsymbol{0}_{p}\\ 1\\ 0\end{pmatrix},\ V\tilde{\mathbf{x}}_{i}=\begin{pmatrix}\boldsymbol{0}_{p}\\ \boldsymbol{0}_{p}\\ 0\\ 1\end{pmatrix},V\tilde{\mathbf{x}}_{n+1}=\boldsymbol{0}_{m},$
Then, for any $i,j\in[n]$ with $i\neq j$, the individual elements of the query-key product are exactly
$(\tilde{X}^{\top}K^{\top}Q\tilde{X})_{j,i}=c\left(\langle\phi(\mathbf{x}_{i}),\mathbf{y}_{j}\rangle+\langle\phi(\mathbf{x}_{j}).\mathbf{y}_{i}\rangle\right).$
By applying Lemma B.4, we find that
$(\tilde{X}^{\top}K^{\top}Q\tilde{X})_{j,i}$ $=2c,\quad\text{if $\mathbf{x}_{i,j}=1$ and $\mathbf{x}_{j,i}=1$};$
$(\tilde{X}^{\top}K^{\top}Q\tilde{X})_{j,i}$ $\leq\frac{3}{2}c,\quad\text{otherwise.}$.
In contrast, $(\tilde{X}^{\top}K^{\top}Q\tilde{X})_{n+1,i}=\frac{7}{4}c$ for any $i$.
Thus, for sufficiently large $c$, all nonzero elements (after rounding) of $\text{sm}(\tilde{X}^{\top}K^{\top}Q\tilde{X})_{\cdot,i}$ belongs to indices $j\in[n]$ if there exist at least one 2-cycle containing node $i$; if not, then $\text{sm}(\tilde{X}^{\top}K^{\top}Q\tilde{X})_{n+1,i}=1$ and all others are zero.
By our choice of value vectors, the $i$th output of the self-attention unit is $\mathbf{e}_{m}$ if there exists a 2-cycle and $\boldsymbol{0}_{m}$ otherwise. ∎
Appendix C Proofs from Section 5
C.1 Proof of Theorem 5.1
See 5.1
Proof.
The main bulk of the proof will use the transformer to prepare the inputs. We will first explain the layout of the construction, and then present it formally. Each input node is represented as a row of the adjacency matrix. We will split the nodes into $n^{1/k}$ sets, where each set contains $n^{1-1/k}$ nodes. The first layer will prepare the adjacency rows so that each token will include only edges of other nodes from the same set. The second layer will combine all the nodes of each set into a separate token. This will use $n^{1/k}$ tokens, where each of them will contain at most $n^{2-2/k}$ edges. The last layer will use each token to represent each possible combination of $k$ such sets. There are at most $\binom{n^{1/k}}{k}\leq n$ such combinations, and each of them contains at most $n^{2-2/k}$ edges. We need an additional $n^{1/k}$ entries for technical reasons to do this embedding into all possible combinations of sets.
We now turn to the formal construction. Assume that the nodes are numbered as $v_{1},\dots,v_{n}$, and denote by $\mathbf{x}_{1},\dots,\mathbf{x}_{n}$ the row of the adjacency matrix corresponding to the nodes. The input to the transformer of the node $v_{i}$ will be $\begin{pmatrix}\mathbf{x}_{i}\\ i\end{pmatrix}\in\mathbb{R}^{n+1}$, where $\mathbf{e}_{i}$ is the $i$-th standard unit vector.
Throughout the proof we assume that $n^{1/k}$ and $n^{1-1/k}$ are integers. Otherwise, replace them by their integral value.
Layer 1:
We begin the construction with an MLP that operates on each token separately. This can be viewed as if we use the self-attention layer to have no effect on the inputs, by setting $V=0$ and using the residual connection. The MLP will implement the following function:
$\mathbb{R}^{n+1}\ni\begin{pmatrix}\mathbf{x}_{i}\\ i\end{pmatrix}\mapsto\begin{pmatrix}\tilde{\mathbf{x}}_{i}\\ \mathbf{w}_{i}\\ \mathbf{z}_{i}\\ i\end{pmatrix}\in\mathbb{R}^{n^{2-2/k}+2n^{1/k}+1}~{}.$
Intuitively, we split the nodes into $n^{1/k}$ sets, each one containing $n^{1-1/k}$ nodes. $\tilde{\mathbf{x}}_{i}$ will include a pruned adjacency row for node $i$ with only edges from its own set. $\mathbf{w}_{i}$ indicates to which set each node belongs to, and $\mathbf{z}_{i}$ indicate on the tokens that will store these sets. We first introduce the vectors $\bar{\mathbf{x}}_{i}\in\mathbb{R}^{n^{1-1/k}}$ that are equal to:
$(\bar{\mathbf{x}}_{i})_{j}=\sum_{r=1}^{n^{1/k}}\mathds{1}((\mathbf{x}_{i})_{(r-1)n^{1-1/k}+j}=1)\cdot\mathds{1}((r-1)n^{1-1/k}+1\leq i\leq rn^{1-1/k})\cdot\mathds{1}((r-1)n^{1-1/k}+1\leq j\leq rn^{1-1/k})$
These vectors can be constructed using a $3$-layer MLP. First, note that this function in our case operates only on integer value inputs, since $i,j$ and all the entries of $\mathbf{x}_{i}$ are integers, hence it is enough to approximate the indicator function up to a uniform error of $\frac{1}{2}$ and it will suffice for our purposes. To this end, we define the function:
$f_{r,s}(z)=\sigma(x-(r-1))-\sigma(x-r)+\sigma(x-(s+1))-\sigma(x-s)~{}.$
Here $\sigma=\max\{0,z\}$ is the ReLU function. If $s\geq r+2$ We get that $f_{r,s}(z)=1$ for $r\leq z\leq s$, and $f_{r,s}(z)=0$ for $z\leq r-1$ or $z\geq s+1$. This shows that the functions inside the indicators can be expressed (for integer valued inputs) using a $2$-layer MLP. Expressing the multiplication of the indicators can be done using another layer:
$g(z_{1},z_{2},z_{3})=\sigma(z_{1}+z_{2}+z_{3}-2)=\mathds{1}(z_{1}=1)\cdot\mathds{1}(z_{2}=1)\cdot\mathds{1}(z_{3}=1)~{},$
where $z_{1},z_{2},z_{3}\in\{0,1\}$. The width of this construction is $O(n)$ since for each of the $n^{1-1/k}$ coordinates of the output we sum $n^{1/k}$ such functions as above.
We define $\tilde{\mathbf{x}}_{i}\in\mathbb{R}^{n^{2-2/k}}$ for $i\equiv j(\text{mod }n^{1-1/k})$ to be equal to $\bar{\mathbf{x}}_{i}$ in the coordinates $(j-1)n^{1-1/k}+1$ until $jn^{1-1/k}$ and all the other coordinates are 00. These vectors will later be summed together across all nodes in the same set, which provides an encoding of all the edges in the set. We also define $\mathbf{w}_{i}=\mathbf{e}_{j}\in\mathbb{R}^{n^{1/k}}$ for $(j-1)n^{1/k}+1\leq i\leq jn^{1/k}$ and $\mathbf{z}_{i}=\mathbf{e}_{i}\in\mathbb{R}^{n^{1/k}}$ for $i=1,\dots,n^{1/k}$ and $\mathbf{z}_{i}=\boldsymbol{0}$ otherwise.
Layer 2:
We define the weights of the second layer of self-attention in the following way:
$K=\begin{pmatrix}\boldsymbol{0}_{n^{2-2/k}\times n^{2-2/k}}&&\\ &\boldsymbol{0}_{n^{1/k}\times n^{1/k}}&&\\ &&I_{n^{1/k}}&\\ &&&0\end{pmatrix},$
$Q=\begin{pmatrix}\boldsymbol{0}_{n^{2-2/k}\times n^{2-2/k}}&&\\ &I_{n^{1/k}}&&\\ &&\boldsymbol{0}_{n^{1/k}\times n^{1/k}}&\\ &&&0\end{pmatrix},$
$V=\begin{pmatrix}n^{1/k}I_{n^{2-2/k}}&\\ &\boldsymbol{0}_{(2n^{1/k}+1)\times(2n^{1/k}+1)}\end{pmatrix}~{}.$
Given two vectors $\begin{pmatrix}\tilde{\mathbf{x}}_{i}\\ \mathbf{w}_{i}\\ \mathbf{z}_{i}\\ i\end{pmatrix},\begin{pmatrix}\tilde{\mathbf{x}}_{j}\\ \mathbf{w}_{j}\\ \mathbf{z}_{j}\\ j\end{pmatrix}$, which are outputs of the previous layer, we have that: $\begin{pmatrix}\tilde{\mathbf{x}}_{i}\\ \mathbf{w}_{i}\\ \mathbf{z}_{i}\\ i\end{pmatrix}^{\top}K^{\top}Q\begin{pmatrix}\tilde{\mathbf{x}}_{j}\\ \mathbf{w}_{j}\\ \mathbf{z}_{j}\\ j\end{pmatrix}=\left\langle\mathbf{z}_{i},\mathbf{w}_{j}\right\rangle$. This shows that the first $n^{1/k}$ tokens, which represent each of the $n^{1/k}$ sets, will attend with a similar weight to every node in their set. After applying the $V$ matrix we use the residual connection only for the positional embedding vectors [^3] It is always possible to use the residual connection to affect only a subset of the coordinates. This can be done by doubling the number of unaffected coordinates, using the $V$ matrix to move the unaffected entries to these new coordinates, and then using a $1$-layer MLP to move the unaffected entries to their previous coordinates (which now include what was added through the residual connection). We omit this construction from here for brevity and since it only changes the embedding dimension by a constant factor. (namely, the last $12n^{1/k}+$ 1 coordinates). Thus, the output of the self-attention layer for the first $n^{1/k}$ tokens encodes all the edges in their set in their first $n^{2-2/k}$ coordinates. This encoding is such that there is $1$ in the $i$-th coordinates if there is an edge between nodes $v_{s}$ and $v_{r}$ in the set for where $r$ and $s$ are the unique integers such that $i\equiv r(\text{mod }n^{1-1/k})$ and $(s-1)n^{2-2/k}+1<i<sn^{2-2/k}$. Thus, the output of the self-attention layer can be written as $\begin{pmatrix}\mathbf{y}_{i}\\ \mathbf{w}_{i}\\ \mathbf{z}_{i}\\ i\end{pmatrix}\in\mathbb{R}^{n^{2-1/k}+2n^{1/k}+1}$, where $\mathbf{y}_{i}$ is either an encoding as described above (for $i\leq n^{1/k}$) or some other vector (for $i\geq n^{1/k}$) for which its exact value will not matter. The vector $\mathbf{w}_{i}$ is a positional embedding that is not needed anymore and will be removed by the MLP, and $\mathbf{z}_{i}=\mathbf{e}_{i}$ for $i\leq n^{1/k}$ and $\mathbf{z}_{i}=\boldsymbol{0}$ otherwise.
We will now construct the MLP of the second layer. First, note that given $n^{1/k}$ sets, the number of all $k$ combinations of such sets is bounded by $\binom{n^{1/k}}{k}\leq n^{k\cdot 1/k}=n$. Denote all possible combinations by $B_{1},\dots,B_{n}$ and let $\mathbf{v}_{1},\dots,\mathbf{v}_{n}\in\mathbb{R}^{n^{1/k}}$ such that $(\mathbf{v}_{i})_{j}=1$ if $B_{i}$ includes the $j$-th set, and 00 otherwise. These vectors encode all the possible combinations of such sets. The MLP will apply the following map:
$\mathbb{R}^{n^{2-1/k}+2n^{1/k}+1}\ni\begin{pmatrix}\mathbf{y}_{i}\\ \mathbf{w}_{i}\\ \mathbf{z}_{i}\\ i\end{pmatrix}\mapsto\begin{pmatrix}\mathbf{y}_{i}\\ \mathbf{v}_{i}\\ \mathbf{z}_{i}\end{pmatrix}\in\mathbb{R}^{n^{2-1/k}+2n^{1/k}+1}~{}.$
This map can be implemented by a $3$-layer MLP. Specifically, the only part coordinates that changes are those of $\mathbf{w}_{i}$ which are replaced by $\mathbf{v}_{i}$. This can be done using the function $f(i)=\sum_{j=1}^{n}\mathds{1}(i=j)\cdot\mathbf{v}_{j}$, and its construction is similar to the construction of the MLP in the previous layer.
Layer 3:
The last self-attention layer will include the following weight matrices:
$K=\begin{pmatrix}\boldsymbol{0}_{n^{2-1/k}\times n^{2-1/k}}&&\\ &I_{n^{1/k}}&\\ &&\boldsymbol{0}_{n^{1/k}\times n^{1/k}}\end{pmatrix},$
$Q=\begin{pmatrix}\boldsymbol{0}_{n^{2-1/k}\times n^{2-1/k}}&&\\ &\boldsymbol{0}_{n^{1/k}\times n^{1/k}}&\\ &&I_{n^{1/k}}\end{pmatrix},$
$V=\begin{pmatrix}kI_{n^{2-1/k}}&\\ &\boldsymbol{0}_{(2n^{1/k}+1)\times(2n^{1/k})}\end{pmatrix}~{}.$
After applying this layer to the outputs of the previous layer, each token $i$ will attend, with similar weight, to all the sets (out of the $n^{1/k}$ sets of nodes) that appear in its positional embedding vector $\mathbf{v}_{i}$. Thus, after applying this layer, The first $n^{2-1/k}$ contain an encoding (as described in the construction of the previous layer) of all the edges in the $i$-th combination of $k$ sets $B_{i}$.
Finally, the MLP will be used to detect whether the given subgraph of size $k$ appears as a subgraph in the input graph (which is an encoding of the edges). The output of the MLP will be $1$ if the subgraph appears and 00 otherwise.
Note that any subgraph of size $k$ must appear in one of those combination of sets. Thus, by summing all the tokens, if their sum is greater than 00 the subgraph of size $k$ appears as a subgraph of $G$.
C.2 Proof of Theorem 5.3
We first define the Eulerian cycle verification problem on multi-graphs.
Consider some directed multi-graph $G=(V,E)$ for $V=\{v_{1},\dots,v_{n}\}$ and $E=\{e_{1},\dots,e_{N}\},$ where each edge is labeled as
$e_{j}=(e_{j,1},e_{j,2},j)\in V\times V\times[N].$
We say that $e_{j}$ is a successor edge of $e_{i}$ if $e_{i,2}=e_{j,1}$. A problem instance also contains a fragmented path, which is expressed as a collection of ordered pairs of edges $P=\{p^{1},\dots,p^{N}\}$ with
$p^{j}=(p^{j}_{1},p^{j}_{2})\in E\times E,$
where $p^{j}_{1}$ and $p^{j}_{2}$ are successive edges (i.e. $p^{j}_{1,2}=p^{j}_{2,1}$). Let $p^{j}$ be a successor path fragment of $p^{i}$ if $p^{i}_{2}=p^{j}_{1}$.
We say that $P$ verifies an Eulerian cycle if
  1. 1.
    every edge in $E$ appears exactly two pairs in $P$; and
  2. 2.
    there exists a permutation over pairs $\sigma:[N]\to[N]$ such that each $p^{\sigma(j+1)}$ is a successor of $p^{\sigma(j)}$ (and $p^{\sigma(0)}$ is a successor of $p^{\sigma(N)}$).
We treat Eulerian cycle detection as a sequential task on adjacency-node tokenization inputs by setting the $i$th embedding to $\phi(v_{i},P_{i})$, where $P_{i}$ encodes all pairs incident to node $v_{i}$, i.e.,
$P_{i}=\{p\in P:p_{1,2}=p_{2,1}=v_{i}\}.$
Now, we prove that—conditional on the hardness of distinguishing one-cycle and two-cycle graphs—no transformer can solve the Eulerian cycle verification problem of degree-$n$ multi-graphs without sufficient width or depth.
See 5.3
Proof.
Consider some transformer $T$ with depth $L$ and embedding dimension $m$ that solves the Eulerian cycle verification problem for any directed multi-graph with $n$ nodes and at most $n^{2}$ edges. We use this to construct a transformer $\mathfrak{T}$ with depth $L+O(1)$ and embedding dimension $O(m+n^{1.1})$ that distinguishes between an undirected444 The one-cycle vs two-cycle conjecture applies only to undirected graphs. cycle graph of size $N$ and two cycles of size $\frac{N}{2}$ for $N=\frac{n^{2}}{2}$. The claim of the theorem follows as an immediate consequence of the one-cycle vs two-cycle conjecture (as stated in Conjecture 13 of Sanford et al. (2024a)).
We prove that a transformer with $O(1)$ layers and embedding dimension $O(n)$ can convert an $N$-node cycle graph instance $\mathfrak{G}=(\mathfrak{V},\mathfrak{E})$ into a multi-graph $G=(V,E)$ with paths $P$ such that $\mathfrak{G}$ is a single cycle if and only if $P$ represents an Eulerian path on $G$. We first define the transformation and then show that it can be implemented by a small transformer.
  • Assume without loss of generality that $\mathfrak{V}=[N]$ and $V=[n]$. Let $\phi_{n}(i)=i\pmod{\frac{n}{2}}$ be a many-to-one mapping from vertices in $\mathfrak{V}$ to half of the vertices in $V$.
  • For each undirected edge $\mathfrak{e}_{i}=\{\mathfrak{v}_{1},\mathfrak{v}_{2}\}\in\mathfrak{E}$, we add two directed edges to $E$:
    $e_{i}=(\phi_{n}(\mathfrak{v}_{1}),\phi_{n}(\mathfrak{v}_{2}),i),\ \text{and}\ e_{-i}=(\phi_{n}(\mathfrak{v}_{2}),\phi_{n}(\mathfrak{v}_{1}),-i).$
    For an arbitrary turnaround edge edge $\mathfrak{e}_{i}=\mathfrak{e}^{*}\in\mathfrak{E}$, we replace $\mathfrak{e}_{i^{*}},\mathfrak{e}_{-i^{*}}$ with two self edges:
    $e_{i}=(\phi_{n}(\mathfrak{v}_{2}),\phi_{n}(\mathfrak{v}_{2}),i),\ \text{and}\ e_{-i}=(\phi_{n}(\mathfrak{v}_{1}),\phi_{n}(\mathfrak{v}_{1}),-i).$
  • For edge $\mathfrak{e}_{i}\in\mathfrak{E}$ as above with unique neighbors $\mathfrak{e}_{j}=\{\mathfrak{v}_{0},\mathfrak{v}_{1}\}$ and $\mathfrak{e}_{k}=\{\mathfrak{v}_{2},\mathfrak{v}_{3}\}$, we add paths $p^{i}=(e_{i},e_{ak})$ and $p^{-i}=(e_{-i},e_{bj})$, where $a$ and $b$ are chosen such that $e_{ak}$ and $e_{bj}$ succeed $e_{i}$ and $e_{-i}$ respectively.
We remark on a few properties of the constructed graph $G$, which satisfies $|V|=n$, and $|E|=|P|=n^{2}$. Any two adjacent edges in $\mathfrak{G}$ create two “successor relationships” between pairs of path segments in $P$. Then, if a cyclic subgraph of $\mathfrak{G}$ does not contain $\mathfrak{e}^{*}$, the path segments in $P$ produced by the edges in the subgraph comprise two disjoint directed cycle paths. On the contrary, if the subgraph contains $\mathfrak{e}^{*}$, then its path segments comprise a single cycle path.
Therefore, if $\mathfrak{G}$ is contains a single cycle of length $N$, then $P$ verifies an Eulerian cycle in $G$ that includes all $n^{2}$ edges Otherwise, if $\mathfrak{G}$ has two cycles of length $2N$, then $P$ represents three cyclic paths, one of length $\frac{n^{2}}{2}$ and two of length $\frac{n^{2}}{4}$. Hence, there is a one-to-one correspondence between the one-cycle vs two-cycle detection problem on $\mathfrak{G}$ and the Eulerian cycle verification problem on $G$ and $P$.
We conclude by outlining the construction of transformer $\mathfrak{T}$ that solves the cycle distinction problem. This can be implemented using elementary constructions or the existing equivalence between transformers and MPC.
  • $\mathfrak{T}$ takes as input a stream of edges, $\mathfrak{e}_{1},\dots\mathfrak{e}_{N}\in\mathfrak{E}$, in no particular order. These are expressed in the edge tokenization, which means that the $i$th input to the transformer is $(\mathfrak{e}_{i,1},\mathfrak{e}_{i,2},i)$. We arbitrarily denote $\mathfrak{e}_{1}=\mathfrak{e}^{*}$ with its positional embedding.
  • In the first attention layer, $\mathfrak{T}$ retrieves the two adjacent edges of each edge embedding.
    It does so with two attention heads. The first encodes $\mathfrak{e}_{i,1}$ as a query vector (which is selected to be nearly orthogonal to those of each of the other $N$ edge embeddings), $\mathfrak{e}_{i,1}+\mathfrak{e}_{i,2}$ as a key vector, and $(\mathfrak{e}_{i,1},\mathfrak{e}_{i,2},i)$ as the value vector. The second does the same with $\mathfrak{e}_{i,2}$. $O(\log n)$ embedding dimension suffices for this association.
    Then, the $i$th output of this layer is processed by an MLP that computes $p^{i}$ and $p^{-i}$.
  • The second attention layer collects all path tokens incident to node $j\in[n]$ (i.e., every $p^{i}\in P_{j}$) in the $j$th embedding. This can be treated as a single communication operation in the Massively Parallel Computation (MPC) model of Karloff et al. (2010), where $N$ machines each send $O(\log n)$ bits of information to $n$ machines, where each machine receives at most $O(n\log n)$ bits. Due to Theorem 1 of Sanford et al. (2024a), this attention layer can complete the routing task with embedding dimension $m=O(n^{1.1})$.
  • Now, the $j$th element computes $\phi(v_{j},P_{j})$ and passes the embedding as input to $T$.
  • If $T$ verifies an Eulerian cycle, let $\mathfrak{T}$ output that $\mathfrak{G}$ is a single-cycle graph.
Therefore, the existence of a transformer $T$ with depth $o(\log n)$ or width $O(N^{2-\epsilon})$ that solves the Eulerian cycle verification problem contradicts the one-cycle vs two-cycle conjecture. This completes the proof. ∎
Appendix D Additional Experiments
Here, we provide additional results of the experiment described in Section 6.1. The results for the 4-Cycle and Triangle Count with 50 and 100 nodes, as well as the connectivity task with 50 nodes, are presented in Figures  4, 5, 6, 7, 8. These experiments present the same trend discussed in Section 6.1, where the training loss and test accuracy and loss are similar, while training and inference times are drastically better for shallow-wide networks.
In Figure 9 we present the critical width chart as described in Section 6.1, for the Triangle Count task.
Figure 4: Training and inference times (a), training loss curves (b), and accuracy curves (c) for the 4-Cycle count task over graphs with $100$ nodes, across transformers with approximately 100k parameters, varying in width and depth. While the loss and accuracies remain consistent, shallow and wide transformers demonstrate significantly faster training and inference times.
Figure 5: Training and inference times (a), training loss curves (b), and accuracy curves (c) for the 4-Cycle count task over graphs with $100$ nodes, across transformers with approximately 100k parameters, varying in width and depth. While the loss and accuracies remain consistent, shallow and wide transformers demonstrate significantly faster training and inference times.
Figure 6: Training and inference times (a), training loss curves (b), and accuracy curves (c) for the 4-Cycle count task over graphs with $50$ nodes, across transformers with approximately 100k parameters, varying in width and depth. While the loss and accuracies remain consistent, shallow and wide transformers demonstrate significantly faster training and inference times.
Figure 7: Training and inference times (a), training loss curves (b), and accuracy curves (c) for the Triangle Count task over graphs with $50$ nodes, across transformers with approximately 100k parameters, varying in width and depth. While the loss and accuracies remain consistent, shallow and wide transformers demonstrate significantly faster training and inference times.
Figure 8: Training and inference times (a), training loss curves (b), and accuracy curves (c) for the Triangle Count task over graphs with $100$ nodes, across transformers with approximately 100k parameters, varying in width and depth. While the loss and accuracies remain consistent, shallow and wide transformers demonstrate significantly faster training and inference times.
Figure 9: Critical width evaluation for the Triangle count Task. The points indicate the critical width at which the model fails to fit the data.
Appendix E Experimental Details
Dataset information
In Section 6.1 we used three synthetic datasets, including connectivity, Triangle Count and 4-Cycle Count. The Triangle Count and 4-Cycle Count were presented in (Chen et al., 2020). Each of these datasets contains $5000$ graphs, and the number of nodes is set according to the configuration, as we tested graphs with increasing numbers of nodes. The counting datasets are generated using Erdős–Rényi graphs with an edge probability of $0.1$.
For the connectivity dataset, to avoid a correlation between connectivity and edge probability as exists in Erdős–Rényi graphs, we generated the datasets using diverse graph distributions, each with multiple distribution parameters. The dataset consists of graphs that are either connected or disconnected, generated using different random graph models to ensure diversity. The Erdős–Rényi model $G(n,p)$ is used, where each edge is included independently with probability $p$, and connected graphs are ensured by choosing $p\geq\frac{\ln n}{n}$, while disconnected graphs use a lower $p$. Random Geometric Graphs (RGGs) are also employed, where nodes are placed randomly in a unit space, and edges are formed if the Euclidean distance is below a certain threshold $r$; connected graphs use a sufficiently high $r$, whereas disconnected graphs are created with a lower $r$. Additionally, Scale-Free networks generated using the Barabási–Albert model are included, where new nodes attach preferentially to high-degree nodes, ensuring connectivity when enough edges per node ($m$) are allowed, while disconnected graphs are produced by limiting interconnections between components. Lastly, the Stochastic Block Model (SBM) is used to generate community-structured graphs, where intra-community connection probabilities ($p_{\text{intra}}$) are set high for connected graphs, and inter-community probabilities ($p_{\text{inter}}$) are set to zero to ensure disconnected graphs. Each type of graph is sampled in equal proportions, shuffled, and split into training, validation, and test sets to maintain class balance.
In Section 6.3 we used three molecular property prediction datasets from Open Graph Benchmark (OGB) (Hu et al., 2020). In ogbg-molhiv, the task is to predict whether a molecule inhibits HIV replication, a binary classification task based on molecular graphs with atom-level features and bond-level edge features. ogbg-bbbp involves predicting blood-brain barrier permeability, a crucial property for drug development, while ogbg-bace focuses on predicting the ability of a molecule to bind to the BACE1 enzyme, associated with Alzheimer’s disease. Dataset statistics are presented in Table 2.
Hyper-Parameters
For all experiments, we use a fixed drouput rate of 0.1 and Relu activations. In Section 6.1 we tuned the learning rate in $\{10^{-4},5\cdot 10^{-5}\}$, batch size in $\{32,64\}$. In Section 6.3 we tuned the learning rate in $\{10^{-3},5\cdot 10^{-3}\}$, number of layers in $\{3,5,6,10,12\}$, hidden dimensions in $\{32,64\}$. We used batch size of size 64.
Edge List tokenization
Tokenization of the graph as a list of edges is done as follows. Assume a graph over $n$ nodes. Each node is represented by a one-hot encoding vector $b_{i}\in\mathbb{R}^{n}$ concatenated with the node input features $x_{i}\in\mathbb{R}^{d}$. Then, each edge is represented by concatenating its node representations. Each edge representation is fed as an independent token to the transformer. As graphs vary in size, we pad each node representation with zeros to match the maximal graph size in the dataset.
Adjacency Rows tokenization
Tokenization of the graph as an adjacency rows is done as follows. Assume a graph over $n$ nodes and adjacency matrix $A$. Each node is associated with a vector of features $x_{i}\in\mathbb{R}^{d}$. We concatenate to each row of $A$ the node’s corresponding feature vector. This results in a vector of size $n+d$ for each node. As graphs vary in size, we pad each node vector with zeros to match the maximal graph size in the dataset. Each such vector is used as an input token to the transformer
Laplacian Eigenvectors tokenization
Tokenization of the graph as Laplacian eigenvectors is done as follows. Assume a graph over $n$ nodes and graph Laplacian $L$. Each node is associated with a vector of features $x_{i}\in\mathbb{R}^{d}$. We compute the eigenvector decomposition of the graph’s Laplacian. Then, each node is associated an eigenvector and an eigenvalue. We concatenate these two for each node, resulting in a vector of size $n+1$. We then concatenate to each of these spectral vectors the node’s corresponding feature vector $x_{i}$. This results in a vector of size $n+d+1$ for each node. As graphs vary in size, we pad each node vector with zeros to match the maximal graph size in the dataset. Each such vector is used as an input token to the transformer
Table 2: Summary statistics of datasets used in Section 6.
Dataset # Graphs Avg # Nodes Avg # Edges # Node Features # Classes
ogbg-molhiv 41,127 25.5 27.5 9 2
ogbg-molbace 1,513 34.1 36.9 9 2
ogbg-molbbbp 2,039 24.1 26.0 9 2
Small and medium size graph commonly used datasets
In the main paper we mentioned that many commonly used graph datasets contain graphs of relatively small size. Therefore in many real-world cases, the embedding dimension of the model is larger than the size of the graph. Here we list multiple such datasets from the Open Graph Benchmark (OGB) (Hu et al., 2020) as well as TUdatasets (Morris et al., 2020). The datasets, including their statistics, are listed in Table 3.
Table 3: Summary of commonly used graph datasets, where the average number of nodes is relatively small
Dataset # Graphs Avg # Nodes Avg # Edges
ogbg-molhiv 41,127 25.5 27.5
ogbg-molbace 1,513 34.1 36.9
ogbg-molbbbp 2,039 24.1 26.0
ogbg-tox21 7,831 18.6 19.3
ogbg-toxcast 8,576 18.8 19.3
ogbg-muv 93,087 24.2 26.3
ogbg-bace 1,513 34.1 36.9
ogbg-bbbp 2,039 24.1 26.0
ogbg-clintox 1,477 26.2 27.9
ogbg-sider 1,427 33.6 35.4
ogbg-esol 1,128 13.3 13.7
ogbg-freesolv 642 8.7 8.4
ogbg-lipo 4,200 27.0 29.5
IMDB-Binary 1000 19 96
IMDB-Multi 1500 13 65
Proteins 1113 39.06 72.82
NCI1 4110 29.87 32.3
Enzymes 600 32.63 62.14

### v2

@@
-Depth-Width Tradeoffs in Algorithmic Reasoning of Graph Tasks with Transformers
+Depth-Width Tradeoffs for Transformers on Graph Tasks
 Abstract
@@
 In particular, they can be used to solve complex algorithmic problems, including graph-based tasks.
-In such algorithmic tasks a key question is what is the minimal size of a transformer that can implement a task.
+In such algorithmic tasks a key question is what is the minimal size of a transformer that can implement the task.
 Recent work has begun to explore this problem for graph-based tasks, showing that for sub-linear embedding dimension (i.e., model width) logarithmic depth suffices.
-However, an open question, which we address here, is what happens if width is allowed to grow linearly.
+However, an open question, which we address here, is what happens if width is allowed to grow linearly, while depth is kept fixed.
 Here we analyze this setting, and provide the surprising result that with linear width, constant depth suffices for solving a host of graph-based problems.
-This suggests that a moderate increase in width can allow much shallower models, which are advantageous in terms of inference time.
+This suggests that a moderate increase in width can allow much shallower models, which are advantageous in terms of inference and train time.
 For other problems, we show that quadratic width is required.
 Our results demonstrate the complex and intriguing landscape of transformer implementations of graph-based algorithms.
-We support our theoretical results with empirical evaluations.
-Machine Learning, ICML
+We empirically investigate these trade-offs between the relative powers of depth and width and find tasks where wider models have the same accuracy as deep models, while having much faster train and inference time due to parallelizable hardware.
 1 Introduction
-The transformer architecture (Vaswani, 2017), which was initially introduced for machine translation (Bahdanau, 2014), has emerged as the state-of-the-art neural network architecture across many fields, including computer vision (Dosovitskiy et al., 2021) and molecular analysis (Jumper et al., 2021).
-In order to explain how transformers emerged as the dominant neural architecture, it is important to understand the fundamental algorithmic primitives that transformers can solve in an efficient manner.
-To that end, this paper addresses the following question.
-Which architectural properties are necessary for transformers to solve graph algorithmic tasks?
-In this work, we consider transformers that compute properties such as connectivity or the existence of a subgraph on some input graphs.
-We use this to develop a hierarchy of algorithmic tasks that proves sharp thresholds in the model width.
-Graphs provide a compelling testbed for understanding transformer reasoning for serveral reasons.
-First, they serve as a natural “algorithmic playground,” encompassing a wide range of well-known problems that span computational classes.
+The transformer architecture (Vaswani, 2017), has emerged as the state-of-the-art neural network architecture across many fields, including language (Brown et al., 2020), computer vision (Dosovitskiy et al., 2021) and molecular analysis (Jumper et al., 2021).
+This success prompts a basic question: which algorithms can be implemented using transformers, and how do specific architecture choices affect this capability.
+Understanding these questions could enable practitioners to design more computationally efficient models without compromising expressivity.
+Two significant architectural elements of transformers are the depth (the number of layers) and width (the dimension of their latent representation).
+Increasing either property increases the expressivity of the underlying architecture, but the qualitative differences between “shallow and wide” and “deep and narrow” architectures are poorly understood.
+Given the widespread use of specialized parallel hardware for training and serving modern language models, increasing the width has a limited effect on model latency, provided sufficient memory.
+A wide range of recent theoretical work (see e.g., Merrill and Sabharwal, 2023b; Liu et al., 2023) explores these fundamental expressivity questions for various benchmark tasks and transformer scaling regimes.
+However, a complete understanding of the interplay between architecture and algorithmic capabilities—and in particular the fine-grained interplay between depth and width—remains elusive.
+In this work, we aim to crystallize the emergent algorithmic capabilities of transformers as a function of model depth and width, and demonstrate occasions when, contrary to conventional wisdom, the benefits of width eclipse those of depth.
+Specifically, we focus on the following question:
+What are the algorithmic capabilities obtained by increasing the width of a transformer?
+We ground this question in the context of graph algorithmic tasks, which provide a compelling testbed for understanding transformer reasoning.
+Graph algorithms serve as a natural “algorithmic playground,” encompassing a wide range of well-known problems that span computational classes.
 Many of these tasks have already been investigated as benchmarks for language models (Fatemi et al., 2023).
-Second, graph neural network (GNN) architectures (Gilmer et al., 2017; Kipf & Welling, 2017; Veličković et al., 2018; Hamilton et al., 2017) are a well-established class of models for graph-structured data, but they have well-known limitations on their expressive power (Xu et al., 2018; Loukas, 2019).
-This raises an important question about whether transformers are a remedy for those structural limitations of GNNs.
-Finally, inference-time reasoning in large language models (LLMs), such as chain-of-thought (Wei et al., 2022) and graph-of-thought (Besta et al., 2024) prompting can often be framed as a graph search over some solution space.
-By investigating transformers through the lens of graph algorithms, we aim to deepen our understanding of their algorithmic capabilities and the role of design choices, such as model width and depth.
-Questions about transformer algorithmic capabilities have attracted considerable attention, and several results that characterize what can and cannot be calculated using transformers of a certain size have emerged (e.g.
-Merrill & Sabharwal, 2023b; Liu et al., 2023).
-However, we are still far from a complete understanding of the interplay between architecture and algorithmic capabilities.
-In this work we focus on this problem in the context of graph algorithms, with emphases on the role of model depth and width and on a particular node-adjacency tokenization scheme.
-Recent works (Sanford et al., 2024b, c) developed trade-offs between the abilities of transformers of different depths to solve fundamental graph problems.
-They identified parallelism as a key aspect that distinguishes transformers, proving surprising benefits of employing a transformer rather than a GNN for graph-based tasks.
-For example, while GNNs require depth $O(n)$ to determine whether an input graph with $n$ vertices is connected, transformers can solve this task using only $O(\log n)$ layers of self-attention.
-However, these results rely on restrictive assumptions about how graphs are represented as inputs to neural models.
-First, they tokenize an input graph as a sequence of discrete edge embeddings, rather than more standard node-based embeddings.
-Second, they focus on the scaling regime where the transformer’s embedding dimension (which we refer to interchangeably as its width) is restricted to be much smaller than the size of the input graph.
-While a small width is certainly desirable, practical transformers have large widths with strict limitations on model depth.
-For example, when learning over graphs, the width of the networks is often much larger than the number of vertices in the graph.
-Table 3 shows a list of commonly used graph datasets, including many molecular datasets, where the average graph size in the data is less than $40$.
-This is while commonly used network width is above $64$.
-In this paper, we seek to understand which graph algorithms can be implemented in realistic transformer scaling regimes, such as when the width grows linearly in the number of tokens and the depth is constant, independent of the graph size.
-Concretely, we consider transformers that take as input graphs with adjacency embeddings for each node and place upper and lower bounds on the minimum width needed to solve several tasks.
-We focus on the constant-depth regime, where we do not allow the number of attention layers to scale as a function of the input size.
-In Section 4, we introduce tasks—such as graph connectivity and fixed-length cycle detection—for which linear width is necessary and sufficient for dense graph inputs.
-Bounded-degree assumptions reduce the width threshold for the cycle detection task.
-Section 5 shows that more complex tasks—such as subgraph counting and Eulerian cycle verification—require super-linear width.
-We visualize the hierarchy over transformer widths induced by our collection of positive and negative results in Figure 1, which ranges from local node-level tasks that can be solved with constant width (such as computing the degree of each node), to arbitrary functions that necessitate a quadratic width scaling.
+Furthermore, previous theoretical works have routinely employed these algorithmic tasks to unveil the capabilities and limitations of graph neural networks (GNNs, Gilmer et al., 2017).
+For example, the theoretical analysis of Loukas (2019) revealed lower bounds on the depth and width necessary to determine connectivity.
+Sanford et al.
+(2024a) employed similar analysis to elucidate such trade-offs for transformers.
+However, their results pertain primarily to the powers of depth and a specific edge-list representation.
+Other scaling regimes—such as transformers with fixed depth and variable width—have yet to be explored in the context of this toolkit.
+For such graph reasoning tasks, we are principally concerned with the optimal transformer size scaling as a function of the graph size.
+Specifically, given a graph with $n$ nodes and an encoding of an algorithmic task, such as determining connectivity or counting the number of triangles, we investigate the width necessary and sufficient to solve the task as a function of $n$.
+This fixed-depth, variable-width setting is arguably more pertinent given the aforementioned benefits of width in computational efficiency.
+Indeed, practical applications of transformers to graphs have width that is often much larger than the number of vertices in the graph.
+[^1] See Table 3 for a list of commonly used graph datasets (including molecular datasets) where the average graph size in the data is less than $40$.
+Our theoretical contribution is a novel representational hierarchy that characterizes this depth-width dependence in transformers.
+Surprisingly, our results show that multiple problems of interest can be solved with fixed depth and linear width with respect to $n$.
+We provide empirical support for the advantage of these shallow and wide models.
 […]
-Figure 1: The width complexity hierarchy of graph tasks for transformers with node-adjacency tokenizations and constant depth.
-Each row visualizes the width regimes where the task is solvable (blue) or hard (red).
-We validate our results empirically in Section 6.
-Specifically, we show that transformers with large embedding dimension and low depth achieve similar accuracy to deeper models, while having faster training and inference time, due to GPU parallelization.
+Figure 1: The width complexity hierarchy of graph tasks for transformers with constant depth. Each row visualizes the width regimes where the task is solvable (blue) or hard (red).
+Our results elucidate a hierarchy concerning the minimum width needed for constant-depth transformers to represent solutions to various graph algorithmic tasks.
+Section˜5 establishes the relevance of the linear scaling regime, where the width of the transformer increases linearly in the number of nodes $n$.
+Specifically, it considers tasks, such as graph connectivity and detection of fixed length cycles, for which linear width is necessary and sufficient for dense graph input.
+Section˜5 demonstrates that more complex tasks, including subgraph counting and Eulerian cycle verification, require super-linear and near-quadratic width respectively.
+The resulting hierarchy over transformer widths induced by our collection of positive and negative results is visualized in Figure˜1, which ranges from local node-level tasks that can be solved with constant width (such as computing the degree of each node), to arbitrary functions that require quadratic width scaling.
+Some of the lower bounds are conditional on the well known one vs.
+two cycles conjecture (see section˜B.1 for a formal statement).
+The empirical results of Section˜6 demonstrate the relevance of our results in practice.
+We consider various graph problems, and experiment with varying the depth and width of the model under a constraint on the overall parameter count.
+Our results demonstrate that shallower and wider models are significantly faster than narrow and deep models.
 2 Related Works
@@
 Transformer with arbitrary depth (Wei et al., 2021; Yun et al., 2020) or arbitrarily many chain-of-thought tokens (Malach, 2023) are known to be universal approximators.
-While these universality results are fundamental, they leave open questions about the ability of transformers to solve tasks efficiently, particularly in terms of parameter count and depth.
-To address this, various theoretical techniques have been employed to develop a more precise understanding of the scaling trade-offs involved.
-For instance, Merrill & Sabharwal (2023a) show that constant-depth transformers with polynomial width can be simulated by $\mathsf{TC^{0}}$ Boolean circuits, implying that they cannot solve problems like graph connectivity.
-These results provide essential context for our paper; while we prove that transformers with a constant number of attention layers and bounded width can solve such tasks (e.g.
-in Theorem 4.1), this does not contradict their bounds, and together, they imply that a sufficiently large depth or width is necessary in the multi-layer perceptrons of our constructions.
-Similarly, Hao et al.
+Informally, these results suggest that any algorithm can be simulated with a sufficiently large transformer.
+They, however, leave open more fine-grained questions about more compact transformer implementations of a given algorithm, particularly in terms of parameter count and depth.
+Various theoretical techniques have been employed to develop a more precise understanding of the relevant scaling trade-offs.
+For instance, Merrill and Sabharwal (2023a) show that constant-depth transformers with polynomial width can be simulated by $\mathsf{TC^{0}}$ Boolean circuits, implying that they cannot solve problems like graph connectivity.
+These results provide essential context for our paper, where we introduce scaling regimes where transformers can and cannot learn such tasks.Similarly, Hao et al.
 (2022) identify formal language classes that can and cannot be recognized by hard-attention transformers.
-On the other hand, deeper transformers or those with sufficiently long chain-of-thought reasoning can simulate finite-state automata (Liu et al., 2023; Li et al., 2024).
+Another way to augment transformer expressivity is to use chain-of-thought reasoning (Wei et al., 2022).
+Indeed, it has been shown that sufficiently long chains reasoning can simulate finite-state automata (Liu et al., 2023; Li et al., 2024).
 A different perspective frames transformers in terms of communication complexity.
@@
 (2024a) extend this analogy to define a transformer complexity hierarchy for graph tasks.
-While they focus on an edge-list tokenization and depth-based hierarchy, our work considers a width-based hierarchy with a node-adjacency tokenization, leading to different complexity insights.
-Graph transformers.
-Recent advancements in graph transformer architectures have significantly expanded the capabilities of Graph Neural Networks (GNNs) (Gilmer et al., 2017) by leveraging the power of attention mechanisms (Veličković et al., 2018; Brody et al., 2022) to facilitate dynamic node aggregation based on neighboring structures without relying on positional encodings.
-Building on this foundation, models like GraphBert (Zhang et al., 2020) incorporated graph-structured node embeddings to enhance representation learning.
-The evolution continued with the development of comprehensive graph transformer models that extend beyond traditional message-passing paradigms.
-For instance  Kreuzer et al.
-(2021) proposed graph transformers that generalize sequence-based transformers Subsequent models, including Spectral Attention Networks Kreuzer et al.
-(2021), introduced novel mechanisms for integrating spectral information and enhancing scalability for larger graphs.
-Additionally, GraphGPS (Rampášek et al., 2023) exemplifies the hybrid approach by combining attention mechanisms with message-passing techniques, thereby harnessing the strengths of both methods.
-There is an abundance of possible embedding for graph transformers (Kreuzer et al., 2021; Dwivedi & Bresson, 2021; Kim et al., 2022; Rampášek et al., 2023; Zhang et al., 2024).
-Our work does not focus on comparing between such embeddings, but rather to study theoretically the effect of the embedding dimension.
-The embedding type itself is mainly a technical detail for proving our theoretical guarantees.
+Graph transformers and GNNs
+Research in transformers and graph neural networks (GNNs) reveals a rich overlap, primarily because both are fundamentally message-passing models.
+Substantial research effort has been devoted to the study of hybrid architectures to combine GNNs’ intrinsic encoding of input graphs and transformers’ empirical successes.
+A variety of approaches to model hybridization have been explored, including the incorporation of graph structure directly into attention layers (Veličković et al., 2018; Brody et al., 2022), as subgraph features in positional encodings (Zhang et al., 2020), and as spectral features of the graph Laplacian matrix (Kreuzer et al., 2021).
+The aim of this work is to understand transformer architectural trade-offs rather than to introduce a novel architecture.
+However, the graph encoding schemes employed in this paper draw upon key insights from this literature.
+We primarily utilize a simple adjacency encoding in order to isolate the focus of our study to the impacts of changing the embedding dimension, but we also investigate some theoretical and empirical trade-offs for spectral Laplacian-based encodings (Appendix˜A).
 3 Problem setting and notations
@@
 We consider the following setting of transformers: the input is a sequence of $N$ tokens $\mathbf{x}_{1},\dots,\mathbf{x}_{N}$ where $\mathbf{x}_{i}\in\mathbb{R}^{d_{\text{in}}}$.
-We denote by $X^{(0)}\in\mathbb{R}^{d_{\text{in}}\times N}$ the matrix where each row $i$ is equal to $\mathbf{x}_{i}$.
+We denote by $X^{(0)}\in\mathbb{R}^{d_{\text{in}}\times N}$ the matrix where the $i$-th column is equal to $\mathbf{x}_{i}$.
 Each layer of the transformer applies a self-attention mechanism on the inputs, and then an MLP.
-We denote by the input to layer $\ell$ as $X^{(\ell-1)}$.
-The self-attention at layer $\ell$ with $H$ heads is parameterized by matrices $K^{(\ell)}_{h},Q_{h}^{(\ell)}\in\mathbb{R}^{m_{\ell-1}\times m_{\ell-1}}$, $V^{(\ell)}_{h}\in\mathbb{R}^{m_{\ell-1}\times m_{\ell-1}}$.
-It is defined as:
-$Z^{(\ell)}=\sum_{h=1}^{H}V^{(\ell)}_{h}X^{(\ell-1)}\text{sm}(X^{(\ell-1)^{\top}}K_{h}^{(\ell)^{\top}}Q^{(\ell)}_{h}X^{(\ell-1)})~{},$
-where sm is row-wise softmax.
-The output is of dimension $Z^{(\ell)}\in\mathbb{R}^{m_{\ell-1}\times N}$.
-After that we use a residual connection, so the output of the self-attention layer is $\tilde{X}^{(\ell)}=Z^{(\ell)}+X^{(\ell-1)}$ Finally, we apply an MLP $\mathcal{N}^{(\ell)}:\mathbb{R}^{m_{\ell-1}}\rightarrow\mathbb{R}^{m_{\ell}}$ with ReLU activations on each token separately (i.e.
+We denote the input to layer $\ell$ by $X^{(\ell-1)}$.
+The self-attention at layer $\ell$ with $H$ heads is parameterized by matrices $K^{(\ell)}_{h},Q_{h}^{(\ell)}\in\mathbb{R}^{m\times m}$, $V^{(\ell)}_{h}\in\mathbb{R}^{m\times m}$.
+[^2] Except for the first layer where $K^{(\ell)}_{h},Q_{h}^{(\ell)}\in\mathbb{R}^{d_{\text{in}}\times d_{\text{in}}}$, $V^{(\ell)}_{h}\in\mathbb{R}^{m\times d_{\text{in}}}$ It is defined as:
+$Z^{(\ell)}=\sum_{h=1}^{H}V^{(\ell)}_{h}X^{(\ell-1)}\textsc{sm}(X^{(\ell-1)^{\top}}K_{h}^{(\ell)^{\top}}Q^{(\ell)}_{h}X^{(\ell-1)})~,$
+where sm is row-wise softmax, and $m$ is the hidden embedding dimension.
+The output is of dimension $Z^{(\ell)}\in\mathbb{R}^{m\times N}$.
+This is followed by a residual connection, so that the output of the self-attention layer is $\tilde{X}^{(\ell)}=Z^{(\ell)}+X^{(\ell-1)}$.
+Finally, we apply an MLP $\mathcal{N}^{(\ell)}:\mathbb{R}^{m}\rightarrow\mathbb{R}^{m}$ with ReLU activations on each token separately (i.e.
 each column of $\tilde{X}^{(\ell)}$).
-The output of the MLP is denoted as $X^{(\ell)}$, which is the input to the next layer.
-We denote by $\sigma:\mathbb{R}\rightarrow\mathbb{R}$ the ReLU activation.
-We also often use positional encodings which are commonly used in standard transformer architectures.
-It is also common to add normalization layers, however it will not be necessary for our results, and these layers can be trivially added in all of our results without changing the output.
-The bit-precision of all our transformers will be $O(\log(n))$, where $n$ is the number of input tokens.
-This is a common assumption in many previous works (Sanford et al., 2024c, b; Merrill & Sabharwal, 2023b, 2024), and is relatively mild.
+The output of the MLP is $X^{(\ell)}$, and is the input to the next layer.
+We let $\sigma:\mathbb{R}\rightarrow\mathbb{R}$ denote the ReLU activation.
+Critically, we assume that the MLP layer can compute arbitrary functions of each individual token embedding $X^{(\ell)}_{i}$.
+While involved, this assumption is common in the theoretical literature (see e.g.
+Sanford et al., 2024b) and enables a theoretical understanding of the modeling limitations imposed by the attention layer.
+However, we note that in all our constructions, the size of the MLP is at most polynomial in the number of nodes in the input graph.
+Our MLP layers further incorporate arbitrary positional encodings of the index $i$ of each token embedding $\tilde{X}_{i}^{(\ell)}$, i.e.
+$\mathcal{N}^{\ell}(\tilde{X}^{(\ell)})=(g(\tilde{X}_{1}^{(\ell)},1),\dots,g(\tilde{X}_{N}^{(\ell)},N)),$ for some $g$.
+This formalism allows standard transformer positional encodings to be implemented in our theoretical model.
+Normalization layers—which are not explicitly accounted for in our transformer definition—can be similarly incorporated by taking advantage of the arbitrary MLP assumption without changing the results.
+The bit-precision of all our transformers will be $O(\log n)$, where $n$ is the number of input tokens.
+This is a common assumption in many previous works Sanford et al.
+(2024c, b); Merrill and Sabharwal (2023b, 2024), and is relatively mild.
+It is also a necessary requirement for representing a number of size $n$, e.g.
+when having positional embeddings for $n$ tokens.
 We will denote $m=\max(m_{0},\dots,m_{L})$ the embedding dimension, and $d_{\text{in}}$ the input dimension.
 Graph Inputs
-Unlike GNNs, transformers are sequential architectures whose network topology does not encode the structure of an input graph.
-It is therefore a significant modeling decision to decide how a graph is best tokenized into a sequence of embeddings that can be proved to a transformer.
-In this work, we focus primarily on the node-adjacency tokenization scheme, where each token corresponds to a node and encodes its incident edges.
-Given a graph $G$ with $n$ nodes, let $A\in\mathbb{R}^{n\times n}$ be its adjacency matrix.
-The $i$th token input $\mathbf{x}_{i}\in\mathbb{R}^{n}$ to the transformer is defined as the $i$th row of $A$.
-Thus, $X\in\mathbb{R}^{n\times n}$.
-There are numerous alternatives, but we restrict our focus to two others for experimental comparisons in Section 6.
-The edge-list tokenization of Sanford et al.
-(2024a) converts the graph into a sequence of discrete edge tokens.
-For a unique identifier assigned to each $v\in V$, this scheme encodes each edge $(v_{i},v_{j})\in E$ as a token $\mathbf{x}=(i,j)\in\mathbb{R}^{2}$.
-A further alternative is the Laplacian eigenvector tokenization, which is introduced and discussed in Appendix A.
-3.2 Graphs and Parallel Computation
+Because the network topology of transformers (unlike GNNs) does not encode the structure of an input graph, graph structure must be encoded explicitly into the input of the transformer.
+Choosing the format of this tokenization scheme for graph inputs is a significant modeling decision.
+The most fundamental aspect of that choice involves deciding between edge-wise and node-wise schemes.
 Sanford et al.
-(2024b, a) introduced a representational equivalence between transformers with $L$ layers and $O(L)$ rounds of massively parallel computation (MPC).
-MPC (Karloff et al., 2010) is a theoretical formalization of the MapReduce (Dean & Ghemawat, 2008) distributed computing protocol.
-Under MPC, a very large number of machines with limited memory alternate between local computation and global communication rounds.
-The equivalence implies that transformers can act as MPC protocols by tokens as machines, local computation as MLP layers, and communication as self-attention.
-Critically, GNNs are not equivalent to MPC in the same sense, and are rather equivalent to the Local or Congest models, depending the sizes of each message between nodes is bounded (see Loukas, 2019).
-The main advantage of transformers over GNNs is the use of parallelism.
-Graph algorithms are well-understood in the MPC framework.
-For instance, a wide range of graph tasks can be organized into depth-based equivalence classes that depend on the tasks’ computational complexity (Nanongkai & Scquizzato, 2022).
-While few unconditional negative results are known in the MPC framework, the one-cycle vs two-cycle task—which asks whether an input graph is either (1) a single cycle graph with $n$ nodes, or (2) a pair of cycle graphs, with $\frac{n}{2}$ nodes each—is widely expected to require either $\Omega(\log n)$ rounds or local memory of $\Omega(n)$.
-By combining these algorithms and conditional lower bounds with the transformer equivalence, we can obtain a numerous size complexity thresholds for transformers.
-However, the results in Sanford et al. (2024b, a) have two limiting assumptions. The first is that the embedding dimension of the transformer
-is strictly sublinear, namely $m=n^{1-\epsilon}$ for some constant $\epsilon\in(0,1)$ independent of $n$.
-The second limitation is that the graph is represented as an edge list, which requires $\Omega(n^{2})$ tokens for dense graphs.
-In this work we go beyond the strict sublinear regime, and focus on the node-adjacency tokenization strategy.
-Remark 3.1.
-When calculating the complexity of solving certain graph tasks, we usually only count the number of transformer layers, while the MLP can be arbitrarily large.
-This corresponds to the MPC model where usually only the distributed computing time is considered, while it is assumed that there is an unlimited local compute time for each machine.
-Intuitively, the number of self-attention layers is equivalent to the number of distributed computation rounds, while the MLP represent the local computation of each machine.
-4 Beyond Sub-linear Embedding Dimension
-In this section we show that having a transformer with linear embedding dimension can easily improve some known lower bounds based on distributed computing.
-We also show new lower bounds that are stricter for transformers with a linear embedding dimension, and a matching upper bound.
-All the proofs for the theorems in this section are in Appendix B.
-4.1 Improving the sub-linear lower bounds
+(2024a) employ an edge-list tokenization that converts the graph into a sequence of discrete edge tokens.
+However, this scheme suffers computationally for large graphs; the quadratic computational bottleneck of self-attention can result in an $\Omega(n^{4})$ runtime for dense graphs with $n$ nodes.
+In contrast, we consider node-wise encodings, where each node corresponds to exactly one input embedding.
+Our primary theoretical results use what is arguably the simplest such representation: the node-adjacency tokenization scheme.
+In this representation, each input embedding directly encodes a node’s edges as an adjacency vector.
+That is, for graph $G$ with $n$ nodes and adjacency matrix $A\in\{0,1\}^{n\times n}$, the $i$th token input $\mathbf{x}_{i}\in\mathbb{R}^{n}$ to the transformer is defined as the $i$th row of $A$.
+An alternative node-wise scheme is the Laplacian eigenvector tokenization (Kreuzer et al., 2021), which captures global graph structure node-wise in a lossy manner with the most significant eigenvectors of the Laplacian matrix.
+We empirically contrast the node-adjacency scheme with both alternatives in Section˜6, and we further explore the representational properties of the Laplacian tokenization in Appendix˜A.
+The adjacency node embedding is not permutation invariant, similar to the edge-list embedding used in Sanford et al.
+(2024a).
+Namely, if the tokens contain positional embeddings, then changing the order of the tokens may also change the output.
+These embeddings are used for technical convenience, especially since spectral embeddings that are permutation invariant are difficult to analyze for combinatorial problems (such as connectivity, subgraph counting, etc.).
+In section˜6 we provide experiments that justify the use of such embedding on real-world datasets.
+4 The Expressive Power of a Linear Embedding Dimension
+In this section we focus on two graph reasoning tasks: the $1$ vs.
+$2$ cycle problem and the cycle detection problem.
+We show that a fixed-depth transformer can solve these tasks with only linear width.
+We also show that these results are tight.
+All the proofs for the theorems in this section are in Appendix˜B.
+Finally, in Sec.
+4.3 we show that for cycle detection in bounded degree graphs, sublinear width suffices.
+Note that the node degree task as appears in fig.˜1 (namely, calculating the degree of each node) can be easily solved with an $O(\log(n))$ embedding size, as shown e.g., in Sanford et al.
+(2024a)
+4.1 A linear width solution for $1$ vs. $2$ cycle detection
 Previous works on the connection between the MPC model and transformers have shown conditional lower bounds for solving certain tasks.
-One such lower bound include the one-cycle versus two-cycles problem.
-Namely, given a graph with $n$ nodes, distinguish whether it is one cycle of length $n$ or two cycles of length $\frac{n}{2}$.
+One such lower bound includes the $1$ vs.
+$2$ cycle problem.
+Namely, determine whether an $n$-node graph is one cycle of length $n$ or two cycles of length $\frac{n}{2}$.
 Conjecture 13 from Sanford et al.
-(2024a) (see also (Ghaffari et al., 2019)) says that the MPC model with $n^{1-\epsilon}$ memory per machine for any $\epsilon\in(0,1)$ cannot distinguish between the two cases, unless the number of MPC rounds is $\Omega(\log(n))$.
-This also suggests that transformers with an embedding dimension of $O(n^{1-\epsilon})$ cannot solve this task.
-We will first show that having a transformer with linear embedding dimension can solve this task:
+(2024a) (see also Ghaffari et al.
+(2019)) states that any MPC protocol with $n^{1-\epsilon}$ memory per machine for any $\epsilon\in(0,1)$ cannot distinguish between the two cases, unless the number of MPC rounds is $\Omega(\log n)$.
+This condition implies that transformers with an embedding dimension of $O(n^{1-\epsilon})$ cannot solve this task.
+For a formal definition, see section˜B.1.
+We demonstrate the tightness of this bound by showing that a transformer with linear embedding dimension can solve this task:
 Theorem 4.1.
 There exists a transformer with $2$ layers of self-attention, and embedding dimension $O(n)$ that solves the $1$ vs. $2$ cycle problem.
-The proof intuition is that since the graph contains only $n$ edges, we can stack all of them into a single token.
+The proof intuition is that the graph contains only $n$ edges, we can stack all of them into a single token.
 After doing so, we can use the MLP to solve it.
-This result shows that these kind of lower bounds for transformers are brittle in the sense that increasing the embedding dimension even slightly breaks them.
+This demonstrates the brittleness of communication-based transformer lower bounds, since even a slight increase in the embedding dimension breaks them.
 With that said, this result is not surprising, since the graph in this task is very sparse.
@@
 Although this is a lossy compression, it still allows to solve the connectivity problem.
-As a consequence, we can use a similar construction as in Theorem 4.1 where we first apply the sketching transformation to each token (i.e.
+As a consequence, we can use a similar construction as in Theorem˜4.1 where we first apply the sketching transformation to each token (i.e.
 row of the adjacency), then embed all the tokens, into a single token and use the MLP to solve the problem using the algorithm from Ahn et al.
 (2012) (Theorem 3.1).
-Our next results will include problems that cannot be solved by simply compressing all the information of the graph into a single token, and would require a more intricate use of the self-attention layers.
-4.2 The power of a linear embedding dimension
-We have shown that transformers with a linear embedding dimension can improve certain lower bounds that apply for sub-linear dimensions.
-In this section we will show another lower bound on the task of detecting small cycles.
-This lower bound is stronger than the one for the one-cycle versus two-cycles in the sense that the depth-embedding dimension trade-off is even more strict.
-Consider the task of detecting $2$-cycles in a directed graph.
-Namely, detecting whether there are nodes $u$ and $v$ with two edges, one directed from $u$ to $v$ and another directed from $v$ to $u$.
-We can prove the following lower bounds on this task:
+Note that this construction works only in high probability rather than deterministically, since the linear sketching requires using a random matrix, which compresses successfully the adjacency matrix only with some high probability.
+Thus, we don’t present it here formally, but rather as a proof sketch.
+Our next results include problems that cannot be solved by simply compressing all the information of the graph into a single token, and require a more intricate use of the self-attention layers.
+4.2 Linear width is necessary and sufficient for $2$-cycle detection
+This section shows that for the problem of cycle detection, linear width is sufficient for solving the task with fixed-depth, and is also necessary.
+We consider the 2-cycle detection problem for directed graphs.
+The task is to find nodes $u$ and $v$ such that the edges $(u,v)$ and $(v,u)$ exist.
+We begin with the lower bound for this case.
 Theorem 4.2.
@@
 Also, assume that the input graphs to $T$ are embedded such that each token is equal to a row of the adjacency matrix.
-Then, if $T$ can detect $2$-cycles on directed graphs we must have that:
-  1. 1.
-    If $T$ has residual connections then $mpHL=\Omega(n)$.
-  2. 2.
-    If $T$ doesn’t have residual connections then $mpH=\Omega(n)$.
-The proof uses a communication complexity argument, and specifically a reduction to set disjointness problem.
-For a formal definition see Appendix B.
-This lower bound is stronger than the lower bounds in Section 4.1 in the sense that they cannot be improved by having logarithmic depth.
+Then, if $T$ can detect $2$-cycles on directed graphs, the following must hold: (1) If $T$ has residual connections then $mpHL=\Omega(n)$; or (2) If $T$ doesn’t have residual connections then $mpH=\Omega(n)$.
+The proof uses a communication complexity argument, and specifically a reduction to the set disjointness problem.
+For a formal definition see Appendix˜B.
+This lower bound is stronger than the lower bounds in Section˜4.1 in the sense that they cannot be circumvented by logarithmic depth.
 For example, assume that the embedding dimension is $m=O(n^{1-\epsilon})$ for some $\epsilon\in(0,1)$, and that $p,H=O(\log(n))$ (which is often the case in practice).
-Then, transformers with residual connection require a depth of $\Omega(n^{\epsilon})$ to solve this task, which is beyond logarithmic, while transformer without residual connections cannot solve it.
-This lower bound is also unconditional, compared to the lower bouds from (Sanford et al., 2024a) which are conditional on the hardness of the one-cycle versus two-cycles problem (Conjecture 13 there).
+Then, transformers with residual connection require a depth of $\Omega(n^{\epsilon})$ to solve this task, which is beyond logarithmic, while transformers without residual connections cannot solve it.
+This lower bound is also unconditional, compared to the lower bounds from Sanford et al.
+(2024a) which are conditional on the hardness of the $1$ vs.
+$2$ cycle problem (Conjecture 13 therein).
 The caveat of this lower bound is that it relies on having the input graph specifically embedded as rows of the adjacency matrix.
@@
 Namely, that depth $L$ transformers with $\Omega(n)$ embedding dimension can calculate the $L$-th power of an adjacency matrix of graphs.
-In particular, the trace of $A^{L}$ provides the number of cycles of size $L$ in the graph (multiplied by an appropriate constant).
+In particular, the trace of $A^{L}$ equals the number of cycles of size $L$ in the graph (multiplied by an appropriate constant).
 Thus, a $2$-layer transformer with linear embedding dimension can already solve the directed $2$-cycle problem.
@@
 There exists an $O(L)$-layer transformer with embedding dimension $m=O(n)$ such that, for any graph embedded as rows of an adjacency matrix $A$, the output of the transformer in the $i$-th token is the $i$-th row of $A^{L}$.
-The constructive proof of this theorem carefully selects key and query parameters to ensure that the output of the softmax matrix equals $A$.
-This enables an inductive argument that encodes $A^{\ell}$ in the value matrix of the $\ell$th layer, in order to compute $A^{\ell+1}$.
-The above theorem shows that having a transformer with linear embedding dimension in the graph size is very strong in terms of expressivity.
+[…]
+The above theorem shows that having a transformer with embedding dimension proportional to the graph size is representationally powerful.
 Namely, $A^{L}_{i,j}$ counts the number of walks of length $L$ form node $i$ to node $j$.
@@
 Although, there are other algorithm (e.g.
-those presented in Section 4.1) that can solve the connectivity task more efficiently.
-4.3 Sublinear embedding dimension for bounded-degree graphs
+those presented in Section˜4.1) that can solve the connectivity task more efficiently.
+4.3 A sublinear width solution for cycle detection in bounded degree graphs
 The results of the previous section establish that for worst-case graph instance, transformers with node-adjacency tokenizations require a linear embedding dimension to solve simple graph reasoning tasks, such as 2-cycle detection.
@@
 For any $n\in\mathbb{N}$ and $d\leq n$, there exists a single-layer transformer with embedding dimension $O(d\log n)$ that detects 2-cycles in any graph with node degree at most $d$.
-This embedding dimensional is optimal up to logarithmic factors.
-The proof reduces the dimensionality of the input adjacency matrix by incorporating vector embeddings from (Sanford et al., 2024c) into the key, query, and value matrices to produce a “sparse attention unit,” whose activations are large when a respective cycle exists.
-5 Super-linear embedding dimension
-In the previous section we have shown the power of having a transformer with an embedding dimension that is linear in the graph size.
-In this section we study transformers with a super-linear dimension.
-Permitting the embedding dimension to $\Omega(n^{2})$ trivializes the question of expressive power in our setting.
-In this regime, the entire graph can be encoded into a single token, which offsets any algorithmic demands to the computationally-unbounded MLP.
-As a result, we restrict this section’s focus to embedding dimensions that are super-linear, but sub-quadratic, in the number of nodes.
-5.1 Detecting and counting subgraphs
-A very basic graph task is to detect or count a given subgraph.
-This task is important in several fields including biology, organic chemistry, graph kernels and more (See (Jin et al., 2020; Jiang et al., 2010; Pope et al., 2018; Shervashidze et al., 2009) and the discussion in Chen et al.
+This embedding dimension is optimal up to logarithmic factors.
+The proof reduces the dimensionality of the input adjacency matrix by incorporating vector embeddings from Sanford et al.
+(2024c) into the key, query, and value matrices to produce a “sparse attention unit,” whose activations are large when a respective cycle exists.
+5 The expressive power of sub-quadratic embedding dimension
+Thus far, we showed that there are graph problems that can be solved with fixed depth and linear width.
+A natural question is if there are problems where larger width is necessary?
+Intuitively, quadratic width should suffice for solving any task, since it can be used to record the entire graph, and then a sufficiently expressive MLP can solve the task.
+But are there problems where quadratic width is necessary.
+In Sec.
+5.1 we show that for the problem of Eulerian cycle detection, quadratic width is necessary.
+This leads to the interesting question of whether there are problems that require super-linear but sub-quadratic width (i.e., between linear and quadratic).
+We offer a first result in this direction by showing in Sec.
+5.2 that such width is sufficient for the problem of sub-graph counting.
+5.1 Necessity of nearly-quadratic width for Eulerian cycle detection
+As mentioned above, quadratic width trivially suffices for solving graph problems.
+It is therefore interesting to understand whether this upper bound is tight for some graph problems.
+We affirmatively answer this question with the Eulerian cycle verification problem.
+Given a graph and a list of “path fragments,” where each fragment consists of a pair of subsequent edges in a path, the goal of the Eulerian cycle verification problem is to determine whether the properly ordered fragments comprise an Eulerian cycle [^3] Recall that a Eulerian cycle is a cycle over the entire graph that uses each edge exactly once..
+We establish the hardness of this problem on multi-graphs under the well-accepted “$1$ vs.
+$2$ cycle” conjecture.
+Theorem 5.1.
+Under Conjecture 2.4 from Sanford et al.
+(2024b), the Eulerian cycle verification problem on multigraphs with self loops cannot be solved by transformers with adjacency matrix inputs if $m=O\left(n^{2-\epsilon}\right)$ for any constant $\epsilon>0$, unless $L=\Omega(\log(n))$.
+The proof relies on a novel reduction to the $1$ vs.
+$2$ cycle problem.
+The difficulty with the reduction is that the one vs.
+two cycles problem consists of a sparse graph with only $n$ edges and nodes, and to show a sub-quadratic lower bound we need a dense graph with $\Omega(n^{2})$ edges and $n$ nodes.
+To this end, we do a random projection of an instance of the one vs.
+two cycles problem with $n^{2}$ nodes to a multigraph with $n$ nodes, where each node represents $n$ nodes in the original graph, but keeping all the edges.
+This allows us to produce a dense multigraph, for which solving the Eulerian cycle verification problem will solve the one vs.
+two cycles problem in the original graph, before the projection.
+5.2 Sub-quadratic solutions to sub-graph counting
+Subgraph detection and counting are central tasks for fields as diverse as biology, organic chemistry, and graph kernels (see Jin et al.
+(2020); Jiang et al.
+(2010); Pope et al.
+(2018); Shervashidze et al.
+(2009); see also the discussion in Chen et al.
 (2020)).
-This task was also studied in the context of GNNs in Chen et al.
-(2020), where it was shown that $k$-IGN can detect subgraphs of size $k$.
-The caveat of this result is that $k$-IGN require $k$-tensor to construct, which is practically infeasible even for small $k$.
-Here we study the ability of transformer to detect and count subgraphs. Our main result is the following:
+In what follows we show that this problem can be solved with transformers with fixed depth, but width that is between linear and quadratic.
 Theorem 5.2.
 Let $k,n\in\mathbb{N}$, and let $G^{\prime}$ be a graph with $k$ nodes.
-There exists a transformer with $O(1)$ self-attention layers and embedding dimension $O\left(n^{2-1/k}\right)$ such that for any graph $G$ of size $n$,
-counts the number of occurrences of $G^{\prime}$ as a subgraph of $G$.
-The crux of the proof is to implement the seminal ”tri-tri-again“ algorithm (Dolev et al., 2012) using transformers.
-The general idea is that given a graph with $n$ nodes, we split the nodes into $n^{1/k}$ disjoint sets, each containing $n^{1-1/k}$ nodes, where $k$ is the size of the subgraph that should be detected or counted.
-For each possible combination of $k$ such sets we use an MLP to count the given subgraph in it.
-There are $\binom{n^{1/k}}{k}\leq n^{k\cdot 1/k}=n$ such combinations of sets, each one contains at most $n^{2-2/k}$ edges.
+There exists a transformer with $O(1)$ self-attention layers and embedding dimension $O\left(n^{2-1/k}\right)$ that, for any graph $G$ of size $n$, counts the number of occurrences of $G^{\prime}$ as a subgraph of $G$.
+In contrast, known GNN constructions (e.g., see Chen et al., 2020) entail higher-order constructions with $k$-IGN models, which require computing practically-infeasible $k$-order tensors.
+These models require a depth of $k$ to recognize subgraph of size $k$, where each layer uses a $k$-tensor instead of standard matrices, which requires $n^{k}$ parameters each.
+Even for $k=3$ or $k=4$ this may be too large for graphs with a few hundreds of nodes.
+Our proof implements the seminal “tri-tri-again” algorithm (Dolev et al., 2012) using transformers.
+Given a graph with $n$ nodes, we partition the nodes into $n^{1/k}$ disjoint sets, each containing $n^{1-1/k}$ nodes.
+For each possible combination of $k$ such sets, we use an MLP to count the given subgraph in it.
+There are $\binom{n^{1/k}}{k}\leq n^{k\cdot 1/k}=n$ such combinations of sets, each one containing at most $n^{2-2/k}$ edges.
 Thus, each token with a large enough embedding dimension can simulate one combination of subsets, cumulating all the relevant edges.
-Note that subgraph detection is a sub-task of counting, whether the number of occurrences is larger than 00.
-Theorem 5.1 can be compared to Theorem 23 in Sanford et al.
+Note that subgraph detection is a sub-task of counting, where the number of occurrences is larger than 0.
+Theorem˜5.2 can be compared to Theorem 23 in Sanford et al.
 (2024a) that provides a construction for counting $k$-cliques using transformers with sub-linear memory and additional blank tokens.
 There, it was shown that it is possible to count $k$-cliques with a transformer of depth $O(\log\log(n))$, however the number of blank tokens is $O(n^{k-1})$ in the worse case.
-By blank tokens we mean additional empty tokens that are appended to the input and used for scratch space as defined in (Sanford et al., 2024a).
+Here, blank tokens refer to empty tokens that are appended to the input and used for scratch space as defined in Sanford et al.
+(2024a).
 In our result, we require a depth of $O(1)$, and the total number of tokens is $n$, while the embedding dimension is super-linear (but sub-quadratic).
@@
 In fact, it is a common conjecture in the literature that no (possibly randomized) algorithm can detect $k$-cliques in time less than $O(n^{2k/3-\epsilon})$ for any $\epsilon>0$ (see Hypothesis 6 in Williams (2018)).
-5.2 Limitations of sub-quadratic memory
-We now turn to the boundary of the possible embedding dimension for which these kinds of questions are interesting.
-A natural question is whether there is a width that suffices for solving all graph problems.
-Intuitively, this width should be n2 as this is what is required for storing the entire graph.
-Here we show that this is indeed an upper bound on needed depth for an arbitrary problem, and that this bound is tight for a particular problem: the Eulerian cycle verification problem.
-Given a graph and a list of “path fragments,” where each fragment consists of a pair of subsequent edges in a path, the goal of the Eulerian cycle verification problem is to determine whether the properly ordered fragments comprise an Eulerian cycle.
-Recall that a Eulerian cycle is a cycle over the entire graph that uses each edge exactly once.
-We show that this problem on multigraphs with self loops cannot be solved (conditioned on the $1$ vs.
-$2$ cycle conjecture) with a transformers unless its embedding dimensions is quadratic in $n$ or its depth is $\log(n)$.
-Theorem 5.3.
-Under Conjecture 2.4 from Sanford et al.
-(2024b), the Eulerian cycle verification problem on multigraphs with self loops cannot be solved by transformers with adjacency matrix inputs if $m=O\left(n^{2-\epsilon}\right)$ for any constant $\epsilon>0$, unless $L=\Omega(\log(n))$.
 Figure 2: Training and inference times (a), training loss curves (b), and accuracy curves (c) for the connectivity task over graphs with $100$ nodes, across transformers with approximately 100k parameters, varying in width and depth.
@@
 6 Experiments
-In this section, we present experiments to support our claims and theoretical results.
+We showcase the effect of width growth in transformers by training a variety of models on synthetic graph algorithmic tasks.
+Section˜6.1 establishes the relative advantages of scaling width over depth by training a family of models with similar parameter counts and variable network topologies; our results show that the shallower and wider models yield the same accuracy as the deeper models, but with faster training and inference.
+Section˜6.2 evaluates the critical width at which the substructure counting task becomes solvable, and shows that it is roughly linear.
+Section˜6.3 demonstrates the trade-offs of each graph encoding scheme (node-adjacency versus edge-list versus Laplacian-eigenvector) on different tasks.
 [^4] Code is provided in the Supplementary Material.
@@
 More experimental details are provided in Appendix E.
-For all experiments in Section 6.1 and Section 6.2, we used the adjacency rows tokenization as described in Section 5.1.
-Details of the implementation are described in Appendix D.
+For all experiments in Section˜6.1 and Section˜6.2, we used the adjacency rows tokenization as described in Section˜5.2.
+Details of the implementation are described in Appendix˜D.
 We considered the tasks of connectivity, triangles count, and 4-cycle count.
@@
 For the connectivity tasks, we generated synthetic graphs, and the label indicates whether the graph is connected or not.
-All the datasets information is described in detail in  Appendix E.
+All the datasets information is described in detail in  Appendix˜E.
 6.1 Empirical trade-offs between width and depth
@@
 We report the averages over $3$ runs with random seeds.
-The hyper-parameters we tuned are provided in Appendix E.
-The results for the connectivity task over $100$ nodes are presented in Figure 2.
-Additional results for the counting tasks and different graph sizes, present the same trends, and are provided in Appendix D due to space limitations.
-As shown in Figure 2 and Figure 2, the training loss and accuracy remain consistent across all depth and width configurations.
-However, Figure 2 reveals that shallow and wide transformers significantly reduce the total training and inference time compared to their deeper and narrower counterparts.
+The hyper-parameters we tuned are provided in Appendix˜E.
+The results for the connectivity task over $100$ nodes are presented in Figure˜2.
+Additional results for the counting tasks and different graph sizes, present the same trends, and are provided in Appendix˜D due to space limitations.
+As shown in Figure˜2 and Figure˜2, the training loss and accuracy remain consistent across all depth and width configurations.
+However, Figure˜2 reveals that shallow and wide transformers significantly reduce the total training and inference time compared to their deeper and narrower counterparts.
 This may be due to the ability of GPUs to parallelize the computations across the width of the same layer, but not across different layers.
-6.2 Critical Width
-In this subsection, we show that the transformer width required to fit the substructure counting tasks increases roughly linearly with the number of nodes, as argued in  Section 5.1.
+6.2 Minimum width for substructure counting
+Figure 3: Critical width evaluation for the 4-Cycle Count task, using a transformer with 1 layer. The points indicate the critical width at which the model fails to fit the data.
+In this subsection, we show that the transformer width required to fit the substructure counting tasks increases sub-quadratically with the number of nodes, as argued in  Section˜5.2.
 The experiments considered graphs with increasing numbers of nodes, ranging from 50 to 400 in increments of 50, and transformer widths varying from 100 to 800 in increments of 100.
@@
 To determine the critical width, we conducted a grid search over each combination of graph size and model width, and selected the model that fitted the data best.
-The hyper-parameters we considered are provided in Appendix E.
+The hyper-parameters we considered are provided in Appendix˜E.
 To isolate the effect of width, we used a transformer model with one layer.
 We used two attention heads to ensure there exists a width for which the model can fit the data for all the evaluated graph sizes.
-The results for the 4-Cycle count task are presented in Figure 3.
-Due to space limitations, the results for the triangle count task are deferred to Appendix D.
-This figure shows that the critical width increases roughly linearly with the graph size.
-Figure 3: Critical width evaluation for the 4-Cycle Count task, using a transformer with 1 layer. The points indicate the critical width at which the model fails to fit the data.
-6.3 Adjacency Rows Graph Tokenization
+The results for the 4-Cycle count task are presented in Figure˜3.
+Due to space limitations, the results for the triangle count task are deferred to Appendix˜D.
+Figure˜3 shows that the critical width increases roughly linearly with the graph size.
+6.3 Comparison of graph encodings
 In this paper, we focus on a graph tokenization where each row of the graph’s adjacency matrix is treated as a token for the model.
@@
 To validate the effectiveness of this tokenization approach in practice, we evaluate it on real graph datasets.
-We compared the adjacency-row representation to the edge-list representation by training a transformer model on three Open Graph Benchmark (OGB) (Hu et al., 2020) datasets: ogbg-molhiv, ogbg-molbbbp, and ogbg-molbace.
+We compared the adjacency-row representation to the edge-list representation by training a transformer model on three Open Graph Benchmark (OGB) Hu et al.
+(2020) datasets: ogbg-molhiv, ogbg-molbbbp, and ogbg-molbace.
 In ogbg-molhiv, the task is to predict whether a molecule inhibits HIV replication, a binary classification task based on molecular graphs with atom-level features and bond-level edge features.
@@
 ogbg-bace focuses on predicting the ability of a molecule to bind to the BACE1 enzyme, associated with Alzheimer’s disease.
-We also evaluated a tokenization using the eigenvectors and eigenvalues of the graph Laplacian, as commonly used in the literature (Dwivedi & Bresson, 2021; Kreuzer et al., 2021).
-More experimental details, including the dataset statistics can be found in Appendix E.
-The results of our evaluation are summarized in Table 1. In all three tasks, the adj-rows representation achieved better ROC-AUC scores than the edge-list representation.
-In two out of the three, it also improved upon the commonly used Laplacian eigenvectors representations.
-The results suggest that the adjacency representation we use in this paper is empirically effective, and should be considered alongside the commonly used Laplacian eigenvector representation.
-Table 1: This table compares the ROC-AUC performance metrics for different graph representations: Edge List, Adjacency Rows, and Laplacian Eigenvectors (LE).
-The results are reported as the average ROC-AUC over three random seeds, along with the corresponding standard deviations.
+We also evaluated a tokenization using the eigenvectors and eigenvalues of the graph Laplacian, as commonly used in the literature (Dwivedi and Bresson, 2021; Kreuzer et al., 2021).
+More experimental details, including the dataset statistics can be found in Appendix˜E.
+Table 1: ROC-AUC performance metrics for different graph representations: Edge List, Adjacency Rows, and Laplacian Eigenvectors (LE), averaged over 3 random seeds.
 Dataset
@@
 LE 68.11±1.52 55.31±4.79 63.61±2.31
+The results of our evaluation are summarized in Table 1.
+In all three tasks, the adj-rows representation achieved better ROC-AUC scores than the edge-list representation.
+In two out of the three, it also improved upon the commonly used Laplacian eigenvectors representations.
+The results suggest that the adjacency representation we use in this paper is empirically effective, and should be considered alongside the commonly used Laplacian eigenvector representation.
 7 Discussion and Future Work
-This paper uses a collection of graph algorithmic tasks—including subgraph detection, connectivity, and Eulerian cycle verification—to demonstrate the powers of width bounded-depth transformers that take as input node adjacency encodings.
+This paper uses a collection of graph algorithmic tasks—including subgraph detection, one vs two cycles, and Eulerian cycle verification—to demonstrate the powers of width bounded-depth transformers that take as input node adjacency encodings.
 These results include sharp theoretical thresholds that demonstrate the trade-offs between constant, linear, quadratic, and intermediate width regimes.
@@
 There are numerous possible extensions of this work.
-One future direction is to study different graph tokenization schemes, beyond the node-adjacency encoding of this work and edge-list encoding of (Sanford et al., 2024a).
+One future direction is to study different graph tokenization schemes, beyond the node-adjacency encoding of this work and edge-list encoding of Sanford et al.
+(2024a).
 A particularly notable alternative is the smallest eigenvectors of the graph Laplacian, presented as a vector of components for each node.
 This spectral embedding is a standard embedding scheme for GNNs, and the techniques and tasks developed in this paper would likely be relevant to proving similar bounds.
-We provide a preliminary exploration of the trade-offs between the intrinsically local characteristics of adjacency-based tokenization schemes and more global spectral approaches in Appendix A.
-Another future direction is to study the optimization and generalization capabilities of transformers to solve graph problems, beyond the expressiveness limitations presented in this work.
+We provide a preliminary exploration of the trade-offs between the intrinsically local characteristics of adjacency-based tokenization schemes and more global spectral approaches in Appendix˜A.
+Another future direction is to study the optimization and generalization capabilities of transformers to solve graph problems, beyond the expressiveness results presented in this work.
 Acknowledgments
@@
 Most simply, the degree of each node can be computed in a sequence-wise manner with node-adjacency tokenization with embedding dimension $m=1$ by simply computing the inner products $\langle\boldsymbol{1}_{n},\mathbf{x}_{i}\rangle$.
-Constructions like Theorems 4.3 and 4.4 further demonstrate the abilities of adjacency encodings to aggregate local structures.
-In contrast, choosing the smallest eigenvectors in the alternative encoding makes it impossible to even compute each node degree without having embedding dimension $m$ growing linearly in the node count $n$222 Consider the task of computing the degree of a particular node of a graph consisting of $\frac{n}{3}$ disconnected linear subgraphs, each with three nodes connected by two edges.
+Constructions like Theorems˜4.3 and 4.4 further demonstrate the abilities of adjacency encodings to aggregate local structures.
+In contrast, choosing the smallest eigenvectors in the alternative encoding makes it impossible to even compute each node degree without having embedding dimension $m$ growing linearly in the node count nn [^5] Consider the task of computing the degree of a particular node of a graph consisting of $\frac{n}{3}$ disconnected linear subgraphs, each with three nodes connected by two edges.
 The zero eigenvalue thus has multiplicity $\frac{n}{3}$, and hence the eigenvectors $\mathbf{v}_{1},\dots,\mathbf{v}_{n/3}$ exist solely as indicators of connected components.
@@
 The entire graph can be encoded in a single token, which can then convert between $A$ and the spectrum of $\mathcal{L}$.
-Appendix B Proofs from Section 4
-B.1 Proof of Theorem 4.1
+Appendix B Proofs from section˜4
+B.1 The $1$ vs. $2$ cycle conjecture
+The most notable conjecture in distributed computing is the $1$ vs.
+$2$ cycle conjecture, which is a common method for providing conditional lower bounds in the MPC model.
+See Definition 2 in Sanford et al.
+[2024a] for a $(\gamma,\delta)$-MPC protocol.
+In simple words, this is a distributed computing protocol, where the input has length $n$, which is distributed to $\Theta(n^{1+\gamma-\delta})$ machines, each having memory of $s:=\Theta(n^{\delta})$.
+Each communication round, each machine calculates an arbitrary local message bounded by size $s$, and all the messages are sent simultaneously.
+The $1$ vs.
+$2$ cycle conjecture is the following:
+Conjecture B.1 (Conjecture 13 in Sanford et al. [2024a]).
+For any $\gamma>0$ and $\delta\in(0,1)$, any $(\gamma,\delta)$-MPC protocol that distinguishes a single length-$n$ cycle from two disjoint length-$n/2$ uses $\Omega(n)$ communication rounds.
+B.2 Proof of Theorem˜4.1
 See 4.1
@@
 This can be thought of as choosing $V=\boldsymbol{0}$ (while $K$ and $Q$ are arbitrary), and using the residual connection so that the tokens remain the same as the input.
-We now use Lemma B.1 to construct a $3$-layer MLP that changes the embedding of each token such that it includes for each node its neighbors.
-The MLP will does not change the positional encoding, this can be done since ReLU network can simulate the identity function by $z\mapsto\sigma(z)-\sigma(-z)$.
-We add another layer to the MLP that maps $\mathbb{R}^{3}\ni\mathbf{v}_{i}\mapsto\mathbf{u}_{i}\in\mathbb{R}^{3n}$, where $(\mathbf{u}_{i})_{3(i-1)+1:3i}=\mathbf{v}_{i}$ and all the other entries of $\mathbf{u}_{i}$ are equal to 00.
+We now use Lemma˜B.2 to construct a $3$-layer MLP that changes the embedding of each token such that it includes for each node its neighbors.
+The MLP does not change the positional encoding, this can be done since ReLU networks can simulate the identity function by $z\mapsto\sigma(z)-\sigma(-z)$, where $\sigma$ represents the ReLU function.
+We add another layer to the MLP that maps $\mathbb{R}^{3}\ni\mathbf{v}_{i}\mapsto\mathbf{u}_{i}\in\mathbb{R}^{3n}$, where $(\mathbf{u}_{i})_{3(i-1)+1:3i}=\mathbf{v}_{i}$ and all the other entries of $\mathbf{u}_{i}$ are equal to 0.
 The second layer of self-attention will have the following matrices: $K=Q=\boldsymbol{0}_{3n\times 3n},V=n\cdot I_{3n}$.
@@
 Thus, the entire information of the graph is embedded in this vector.
-Finally, we use the MLP to determine whether the input graph, whose embedded as a list of edges, is connected.
-This can be done by an MLP since it has the universal approximation property (Cybenko, 1989; Leshno et al., 1993).
+Finally, we use the MLP to determine whether the input graph, whose edges are embedded as a list of edges, is connected.
+This can be done by an MLP since it has the universal approximation property Cybenko [1989], Leshno et al.
+[1993].
 Although we don’t specify the exact size of this MLP, it can be bounded since there are efficient deterministic algorithms for determining connectivity.
 These algorithms can be simulated using ReLU networks.
-Note that the output of the connectivity problem is either 00 or $1$, thus it is enough to approximate a solution of this task up to a constant error (say, of $\frac{1}{4}$), and then use another layer to threshold over the answer.
+Note that the output of the connectivity problem is either 0 or $1$, thus it is enough to approximate a solution of this task up to a constant error (say, of $\frac{1}{4}$), and then use another layer to threshold over the answer.
 Lemma B.2.
@@
 This is a linear combination of indicators, where each indicator can be implemented by the function $f(z)=\sigma(x)-\sigma(x-1)$.
-Note that the input to the MLP is either 00 or $1$ in each coordinate, thus the output of this function will be $\begin{pmatrix}i+j\\ i^{2}+j^{2}\end{pmatrix}$.
+Note that the input to the MLP is either 0 or $1$ in each coordinate, thus the output of this function will be $\begin{pmatrix}i+j\\ i^{2}+j^{2}\end{pmatrix}$.
 We have that $i+j$ and $i^{2}+j^{2}$ determine the values of $i$ and $j$.
@@
 Since there are $O(n^{2})$ different possible values, we can construct a network with $2$-layers and $O(n^{2})$ width that outputs $\begin{pmatrix}i\\ j\end{pmatrix}$ up to changing the order of $i$ and $j$.
-B.2 Proof of Theorem 4.2
+B.3 Proof of Theorem˜4.2
 See 4.2
 Proof.
-Our proof relies on a communication complexity lower bound for the set disjointness problem, and is similar to the arguments from (Sanford et al., 2024c; Yehudai et al., 2024).
+Our proof relies on a communication complexity lower bound for the set disjointness problem, and is similar to the arguments from Sanford et al.
+[2024c], Yehudai et al.
+[2024].
 The lower bound for communication complexity is the following: Alice and Bob are given inputs $a,b\in\{0,1\}^{s}$ respectively, and their goal is to find $\max a_{i}b_{i}$ by sending single bit-messages to each other in a sequence of communication rounds.
@@
 Hence, there is no need for more communication rounds, and Alice can finish the rest of the calculations of the transformers using the output tokens of the first layer.
-By the equivalence between the set disjointness and the directed $2$-cycle that was described above, Alice returns $1$ iff the inputs $\max_{i}a_{i}b_{i}=1$, and 00 otherwise.
+By the equivalence between the set disjointness and the directed $2$-cycle that was described above, Alice returns $1$ iff the inputs $\max_{i}a_{i}b_{i}=1$, and 0 otherwise.
 The total number of bits transmitted in this protocol in the case there are residual connection is $O(nmpHL)$, since there are $O(nmpH)$ bits transferred in each layer.
@@
 ∎
-B.2.1 An extension to bounded degree graphs
-In order to prove the optimality result in Theorem 4.4 for the task of determining the existence of a 2-cycle in bounded-degree graphs, we state the following theorem.
+B.3.1 An extension to bounded degree graphs
+In order to prove the optimality result in Theorem˜4.4 for the task of determining the existence of a 2-cycle in bounded-degree graphs, we state the following theorem.
 Theorem B.3.
@@
 Proof.
-The proof is nearly identical to that of Theorem 4.2, except that we alter the reduction to ensure that that the graph possessed by Alice and Bob is of degree $d$.
+The proof is nearly identical to that of Theorem˜4.2, except that we alter the reduction to ensure that that the graph possessed by Alice and Bob is of degree $d$.
 As before, we design a graph $G=(V,E)$ with vertices partitioned into two sets $V_{1},V_{2}$ satisfying $|V_{1}|=|V_{2}|=n$.
@@
 ∎
-B.3 Proof of Theorem 4.3
+B.4 Proof of Theorem˜4.3
 We will need the following lemma for the proof.
@@
 We first have that $X^{\top}K^{\top}QX=A$.
-Since all the values of $A$ are either 00 or $1$, for a sufficiently large $c>0$, the softmax behave similarly to the hardmax and we get that:
-$VX\text{sm}(A)=\begin{pmatrix}A^{2}\cdot\text{deg}(A)^{-1}\\ A\cdot\text{deg}(A)^{-1}\\ \boldsymbol{0}_{d\times d}\end{pmatrix}~{},$
+Since all the values of $A$ are either 0 or $1$, for a sufficiently large $c>0$, the softmax behave similarly to the hardmax and we get that:
+$VX\textsc{sm}(A)=\begin{pmatrix}A^{2}\cdot\text{deg}(A)^{-1}\\ A\cdot\text{deg}(A)^{-1}\\ \boldsymbol{0}_{d\times d}\end{pmatrix}~$,
 where $\text{deg}(A)$ is a diagonal matrix, where its $i$-th diagonal entry is equal to the degree of node $i$.
@@
 Constructing such an MLP can be done by calculating the degree of each token from $A\cdot\text{deg}(A)^{-1}$ and multiplying the first $d$ coordinates of each token by this degree.
-This can be done since the entries of this matrix is either 00 or the inverse of the degree of node $i$, thus it requires only inverting an integer between $1$ and $n$.
-By Lemma B.3 this can be done by a $2$-layer MLP with width $n$.
+This can be done since the entries of this matrix is either 0 or the inverse of the degree of node $i$, thus it requires only inverting an integer between $1$ and $n$.
+By Lemma˜B.4 this can be done by a $2$-layer MLP with width $n$.
 This finishes the construction for calculating $A^{2}$.
@@
 ∎
-B.4 Proof of Theorem 4.4
+B.5 Proof of Theorem˜4.4
 See 4.4
-The optimality result is proved using the same methodology as Theorem 4.2. It is stated and proved formally as Theorem B.2.
-The proof of the construction adapts an argument from Theorem 2 of (Sanford et al., 2024c), which shows that a sparse averaging task can be solved with bounded-width transformers.
-We make use of the following fact, which is a consequence of the Restricted Isometry Property analysis of (Candes & Tao, 2005; Mendelson et al., 2005).
+The optimality result is proved using the same methodology as Theorem˜4.2. It is stated and proved formally as Theorem˜B.3.
+The proof of the construction adapts an argument from Theorem 2 of Sanford et al.
+[2024c], which shows that a sparse averaging task can be solved with bounded-width transformers.
+We make use of the following fact, which is a consequence of the Restricted Isometry Property analysis of Candes and Tao [2005], Mendelson et al.
+[2005].
 Lemma B.5.
@@
 $\langle\phi(\mathbf{x}),\mathbf{y}_{i}\rangle$ $\leq\frac{1}{2},\quad\text{if $\mathbf{x}_{i}=0$.}$
-We use this fact to prove Theorem 4.4.
+We use this fact to prove Theorem˜4.4.
 Proof.
@@
 $(\tilde{X}^{\top}K^{\top}Q\tilde{X})_{j,i}=c\left(\langle\phi(\mathbf{x}_{i}),\mathbf{y}_{j}\rangle+\langle\phi(\mathbf{x}_{j}).\mathbf{y}_{i}\rangle\right).$
-By applying Lemma B.4, we find that
+By applying Lemma˜B.5, we find that
 $(\tilde{X}^{\top}K^{\top}Q\tilde{X})_{j,i}$ $=2c,\quad\text{if $\mathbf{x}_{i,j}=1$ and $\mathbf{x}_{j,i}=1$};$
-$(\tilde{X}^{\top}K^{\top}Q\tilde{X})_{j,i}$ $\leq\frac{3}{2}c,\quad\text{otherwise.}$.
+$(\tilde{X}^{\top}K^{\top}Q\tilde{X})_{j,i}$ $\leq\frac{3}{2}c,\quad\text{otherwise.}$
 In contrast, $(\tilde{X}^{\top}K^{\top}Q\tilde{X})_{n+1,i}=\frac{7}{4}c$ for any $i$.
-Thus, for sufficiently large $c$, all nonzero elements (after rounding) of $\text{sm}(\tilde{X}^{\top}K^{\top}Q\tilde{X})_{\cdot,i}$ belongs to indices $j\in[n]$ if there exist at least one 2-cycle containing node $i$; if not, then $\text{sm}(\tilde{X}^{\top}K^{\top}Q\tilde{X})_{n+1,i}=1$ and all others are zero.
+Thus, for sufficiently large $c$, all nonzero elements (after rounding) of $\textsc{sm}(\tilde{X}^{\top}K^{\top}Q\tilde{X})_{\cdot,i}$ belongs to indices $j\in[n]$ if there exist at least one 2-cycle containing node $i$; if not, then $\textsc{sm}(\tilde{X}^{\top}K^{\top}Q\tilde{X})_{n+1,i}=1$ and all others are zero.
 By our choice of value vectors, the $i$th output of the self-attention unit is $\mathbf{e}_{m}$ if there exists a 2-cycle and $\boldsymbol{0}_{m}$ otherwise. ∎
-Appendix C Proofs from Section 5
-C.1 Proof of Theorem 5.1
-See 5.1
+Appendix C Proofs from Section˜5
+C.1 Proof of Theorem˜5.2
+See 5.2
 Proof.
@@
 where $z_{1},z_{2},z_{3}\in\{0,1\}$. The width of this construction is $O(n)$ since for each of the $n^{1-1/k}$ coordinates of the output we sum $n^{1/k}$ such functions as above.
-We define $\tilde{\mathbf{x}}_{i}\in\mathbb{R}^{n^{2-2/k}}$ for $i\equiv j(\text{mod }n^{1-1/k})$ to be equal to $\bar{\mathbf{x}}_{i}$ in the coordinates $(j-1)n^{1-1/k}+1$ until $jn^{1-1/k}$ and all the other coordinates are 00.
+We define $\tilde{\mathbf{x}}_{i}\in\mathbb{R}^{n^{2-2/k}}$ for $i\equiv j(\text{mod }n^{1-1/k})$ to be equal to $\bar{\mathbf{x}}_{i}$ in the coordinates $(j-1)n^{1-1/k}+1$ until $jn^{1-1/k}$ and all the other coordinates are 0.
 These vectors will later be summed together across all nodes in the same set, which provides an encoding of all the edges in the set.
@@
 We omit this construction from here for brevity and since it only changes the embedding dimension by a constant factor.
-(namely, the last $12n^{1/k}+$ 1 coordinates).
+(namely, the last $2n^{1/k}+1$ coordinates).
 Thus, the output of the self-attention layer for the first $n^{1/k}$ tokens encodes all the edges in their set in their first $n^{2-2/k}$ coordinates.
@@
 First, note that given $n^{1/k}$ sets, the number of all $k$ combinations of such sets is bounded by $\binom{n^{1/k}}{k}\leq n^{k\cdot 1/k}=n$.
-Denote all possible combinations by $B_{1},\dots,B_{n}$ and let $\mathbf{v}_{1},\dots,\mathbf{v}_{n}\in\mathbb{R}^{n^{1/k}}$ such that $(\mathbf{v}_{i})_{j}=1$ if $B_{i}$ includes the $j$-th set, and 00 otherwise.
+Denote all possible combinations by $B_{1},\dots,B_{n}$ and let $\mathbf{v}_{1},\dots,\mathbf{v}_{n}\in\mathbb{R}^{n^{1/k}}$ such that $(\mathbf{v}_{i})_{j}=1$ if $B_{i}$ includes the $j$-th set, and 0 otherwise.
 These vectors encode all the possible combinations of such sets.
@@
 Finally, the MLP will be used to detect whether the given subgraph of size $k$ appears as a subgraph in the input graph (which is an encoding of the edges).
-The output of the MLP will be $1$ if the subgraph appears and 00 otherwise.
+The output of the MLP will be $1$ if the subgraph appears and 0 otherwise.
 Note that any subgraph of size $k$ must appear in one of those combination of sets.
-Thus, by summing all the tokens, if their sum is greater than 00 the subgraph of size $k$ appears as a subgraph of $G$.
-C.2 Proof of Theorem 5.3
+Thus, by summing all the tokens, if their sum is greater than 0 the subgraph of size $k$ appears as a subgraph of $G$.
+C.2 Proof of Theorem˜5.1
 We first define the Eulerian cycle verification problem on multi-graphs.
@@
   1. 1.
-    every edge in $E$ appears exactly two pairs in $P$; and
+    every edge in $E$ appears in exactly two pairs in $P$; and
   2. 2.
@@
 Now, we prove that—conditional on the hardness of distinguishing one-cycle and two-cycle graphs—no transformer can solve the Eulerian cycle verification problem of degree-$n$ multi-graphs without sufficient width or depth.
-See 5.3
+See 5.1
 Proof.
 Consider some transformer $T$ with depth $L$ and embedding dimension $m$ that solves the Eulerian cycle verification problem for any directed multi-graph with $n$ nodes and at most $n^{2}$ edges.
-We use this to construct a transformer $\mathfrak{T}$ with depth $L+O(1)$ and embedding dimension $O(m+n^{1.1})$ that distinguishes between an undirected444 The one-cycle vs two-cycle conjecture applies only to undirected graphs.
+We use this to construct a transformer $\mathfrak{T}$ with depth $L+O(1)$ and embedding dimension $O(m+n^{1.1})$ that distinguishes between an undirected [^7] The $1$ vs.
+$2$ cycle conjecture applies only to undirected graphs.
 cycle graph of size $N$ and two cycles of size $\frac{N}{2}$ for $N=\frac{n^{2}}{2}$.
-The claim of the theorem follows as an immediate consequence of the one-cycle vs two-cycle conjecture (as stated in Conjecture 13 of Sanford et al.
-(2024a)).
+The claim of the theorem follows as an immediate consequence of the $1$ vs.
+$2$ cycle conjecture (as stated in Conjecture 13 of Sanford et al.
+[2024a]).
 We prove that a transformer with $O(1)$ layers and embedding dimension $O(n)$ can convert an $N$-node cycle graph instance $\mathfrak{G}=(\mathfrak{V},\mathfrak{E})$ into a multi-graph $G=(V,E)$ with paths $P$ such that $\mathfrak{G}$ is a single cycle if and only if $P$ represents an Eulerian path on $G$.
@@
 Therefore, if $\mathfrak{G}$ is contains a single cycle of length $N$, then $P$ verifies an Eulerian cycle in $G$ that includes all $n^{2}$ edges Otherwise, if $\mathfrak{G}$ has two cycles of length $2N$, then $P$ represents three cyclic paths, one of length $\frac{n^{2}}{2}$ and two of length $\frac{n^{2}}{4}$.
-Hence, there is a one-to-one correspondence between the one-cycle vs two-cycle detection problem on $\mathfrak{G}$ and the Eulerian cycle verification problem on $G$ and $P$.
+Hence, there is a one-to-one correspondence between the $1$ vs.
+$2$ cycle detection problem on $\mathfrak{G}$ and the Eulerian cycle verification problem on $G$ and $P$.
 We conclude by outlining the construction of transformer $\mathfrak{T}$ that solves the cycle distinction problem.
@@
 This can be treated as a single communication operation in the Massively Parallel Computation (MPC) model of Karloff et al.
-(2010), where $N$ machines each send $O(\log n)$ bits of information to $n$ machines, where each machine receives at most $O(n\log n)$ bits.
+[2010], where $N$ machines each send $O(\log n)$ bits of information to $n$ machines, where each machine receives at most $O(n\log n)$ bits.
 Due to Theorem 1 of Sanford et al.
-(2024a), this attention layer can complete the routing task with embedding dimension $m=O(n^{1.1})$.
+[2024a], this attention layer can complete the routing task with embedding dimension $m=O(n^{1.1})$.
   • Now, the $j$th element computes $\phi(v_{j},P_{j})$ and passes the embedding as input to $T$.
   • If $T$ verifies an Eulerian cycle, let $\mathfrak{T}$ output that $\mathfrak{G}$ is a single-cycle graph.
-Therefore, the existence of a transformer $T$ with depth $o(\log n)$ or width $O(N^{2-\epsilon})$ that solves the Eulerian cycle verification problem contradicts the one-cycle vs two-cycle conjecture.
+Therefore, the existence of a transformer $T$ with depth $o(\log n)$ or width $O(N^{2-\epsilon})$ that solves the Eulerian cycle verification problem contradicts the $1$ vs.
+$2$ cycle conjecture.
 This completes the proof.
@@
 Appendix D Additional Experiments
-Here, we provide additional results of the experiment described in Section 6.1.
+Here, we provide additional results of the experiment described in Section˜6.1.
 The results for the 4-Cycle and Triangle Count with 50 and 100 nodes, as well as the connectivity task with 50 nodes, are presented in Figures  4, 5, 6, 7, 8.
-These experiments present the same trend discussed in Section 6.1, where the training loss and test accuracy and loss are similar, while training and inference times are drastically better for shallow-wide networks.
-In Figure 9 we present the critical width chart as described in Section 6.1, for the Triangle Count task.
+These experiments present the same trend discussed in Section˜6.1, where the training loss and test accuracy and loss are similar, while training and inference times are drastically better for shallow-wide networks.
+In Figure˜9 we present the critical width chart as described in Section˜6.1, for the Triangle Count task.
 Figure 4: Training and inference times (a), training loss curves (b), and accuracy curves (c) for the 4-Cycle count task over graphs with $100$ nodes, across transformers with approximately 100k parameters, varying in width and depth.
@@
 Dataset information
-In Section 6.1 we used three synthetic datasets, including connectivity, Triangle Count and 4-Cycle Count.
-The Triangle Count and 4-Cycle Count were presented in (Chen et al., 2020).
+In Section˜6.1 we used three synthetic datasets, including connectivity, Triangle Count and 4-Cycle Count.
+The Triangle Count and 4-Cycle Count were presented in Chen et al.
+[2020].
 Each of these datasets contains $5000$ graphs, and the number of nodes is set according to the configuration, as we tested graphs with increasing numbers of nodes.
@@
 Each type of graph is sampled in equal proportions, shuffled, and split into training, validation, and test sets to maintain class balance.
-In Section 6.3 we used three molecular property prediction datasets from Open Graph Benchmark (OGB) (Hu et al., 2020).
+In Section˜6.3 we used three molecular property prediction datasets from Open Graph Benchmark (OGB) Hu et al.
+[2020].
 In ogbg-molhiv, the task is to predict whether a molecule inhibits HIV replication, a binary classification task based on molecular graphs with atom-level features and bond-level edge features.
 ogbg-bbbp involves predicting blood-brain barrier permeability, a crucial property for drug development, while ogbg-bace focuses on predicting the ability of a molecule to bind to the BACE1 enzyme, associated with Alzheimer’s disease.
-Dataset statistics are presented in Table 2.
+Dataset statistics are presented in Table˜2.
 Hyper-Parameters
 For all experiments, we use a fixed drouput rate of 0.1 and Relu activations.
-In Section 6.1 we tuned the learning rate in $\{10^{-4},5\cdot 10^{-5}\}$, batch size in $\{32,64\}$.
-In Section 6.3 we tuned the learning rate in $\{10^{-3},5\cdot 10^{-3}\}$, number of layers in $\{3,5,6,10,12\}$, hidden dimensions in $\{32,64\}$.
+In Section˜6.1 we tuned the learning rate in $\{10^{-4},5\cdot 10^{-5}\}$, batch size in $\{32,64\}$.
+In Section˜6.3 we tuned the learning rate in $\{10^{-3},5\cdot 10^{-3}\}$, number of layers in $\{3,5,6,10,12\}$, hidden dimensions in $\{32,64\}$.
 We used batch size of size 64.
@@
 Each such vector is used as an input token to the transformer
-Table 2: Summary statistics of datasets used in Section 6.
+Table 2: Summary statistics of datasets used in Section˜6.
 Dataset # Graphs Avg # Nodes Avg # Edges # Node Features # Classes
@@
 Therefore in many real-world cases, the embedding dimension of the model is larger than the size of the graph.
-Here we list multiple such datasets from the Open Graph Benchmark (OGB) (Hu et al., 2020) as well as TUdatasets (Morris et al., 2020).
-The datasets, including their statistics, are listed in Table 3.
+Here we list multiple such datasets from the Open Graph Benchmark (OGB) Hu et al.
+[2020] as well as TUdatasets Morris et al.
+[2020].
+The datasets, including their statistics, are listed in Table˜3.
 Table 3: Summary of commonly used graph datasets, where the average number of nodes is relatively small
@@
 Enzymes 600 32.63 62.14
+NeurIPS Paper Checklist
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4. Justification:
+  5. Guidelines:
+    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
+    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
+A No or NA answer to this question will not be perceived well by the reviewers.
+    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
+    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
+  6. 2.
+    Limitations
+  7. Question: Does the paper discuss the limitations of the work performed by the authors?
+  8. Answer: [Yes]
+  9. Justification:
+  10. Guidelines:
+    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
+    • The authors are encouraged to create a separate "Limitations" section in their paper.
+    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
+The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
+    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
+In general, empirical results often depend on implicit assumptions, which should be articulated.
+    • The authors should reflect on the factors that influence the performance of the approach.
+For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
+Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
+    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
+    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
+    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
+The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
+Reviewers will be specifically instructed to not penalize honesty concerning limitations.
+  11. 3.
+    Theory assumptions and proofs
+  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  13. Answer: [Yes]
+  14. Justification:
+  15. Guidelines:
+    • The answer NA means that the paper does not include theoretical results.
+    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
+    • All assumptions should be clearly stated or referenced in the statement of any theorems.
+    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
+    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
+    • Theorems and Lemmas that the proof relies upon should be properly referenced.
+  16. 4.
+    Experimental result reproducibility
+  17.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  18. Answer: [Yes]
+  19. Justification:
+  20. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
+    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
+    • Depending on the contribution, reproducibility can be accomplished in various ways.
+For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
+In general.
+releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
+    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
+For example
+      1. (a)
+        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
+      2. (b)
+        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
+      3. (c)
+        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
+      4. (d)
+        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
+In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
+  21. 5.
+    Open access to data and code
+  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  23. Answer: [Yes]
+  24. Justification: We will provide in the final version a Github project with the code.
+  25. Guidelines:
+    • The answer NA means that paper does not include experiments requiring code.
+    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
+Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
+    • The instructions should contain the exact command and environment needed to run to reproduce the results.
+See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
+    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
+If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
+    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
+    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
+  26. 6.
+    Experimental setting/details
+  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+  28. Answer: [Yes]
+  29. Justification:
+  30. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
+    • The full details can be provided either with the code, in appendix, or as supplemental material.
+  31. 7.
+    Experiment statistical significance
+  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  33. Answer: [Yes]
+  34. Justification:
+  35. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
+    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
+    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
+    • The assumptions made should be given (e.g., Normally distributed errors).
+    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
+    • It is OK to report 1-sigma error bars, but one should state it.
+The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
+    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
+    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
+  36. 8.
+    Experiments compute resources
+  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  38. Answer: [Yes]
+  39. Justification:
+  40. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
+    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
+    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
+  41. 9.
+    Code of ethics
+  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  43. Answer: [Yes]
+  44. Justification:
+  45. Guidelines:
+    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
+    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
+    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
+  46. 10.
+    Broader impacts
+  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  48. Answer: [N/A]
+  49. Justification:
+  50. Guidelines:
+    • The answer NA means that there is no societal impact of the work performed.
+    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
+    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
+    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
+However, if there is a direct path to any negative applications, the authors should point it out.
+For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
+On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
+    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
+    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
+  51. 11.
+    Safeguards
+  52.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+  53. Answer: [N/A]
+  54. Justification:
+  55. Guidelines:
+    • The answer NA means that the paper poses no such risks.
+    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
+    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
+    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
+  56. 12.
+    Licenses for existing assets
+  57.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  58. Answer: [Yes]
+  59. Justification:
+  60. Guidelines:
+    • The answer NA means that the paper does not use existing assets.
+    • The authors should cite the original paper that produced the code package or dataset.
+    • The authors should state which version of the asset is used and, if possible, include a URL.
+    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
+    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
+    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
+For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
+Their licensing guide can help determine the license of a dataset.
+    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
+    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
+  61. 13.
+    New assets
+  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  63. Answer: [N/A]
+  64. Justification:
+  65. Guidelines:
+    • The answer NA means that the paper does not release new assets.
+    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
+    • The paper should discuss whether and how consent was obtained from people whose asset is used.
+    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
+  66. 14.
+    Crowdsourcing and research with human subjects
+  67.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  68. Answer: [N/A]
+  69. Justification:
+  70. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
+    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
+  71. 15.
+    Institutional review board (IRB) approvals or equivalent for research with human subjects
+  72.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  73. Answer: [N/A]
+  74. Justification:
+  75. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
+If you obtained IRB approval, you should clearly state this in the paper.
+    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
+    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
+  76. 16.
+    Declaration of LLM usage
+  77.
+Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
+Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
+  78. Answer: [N/A]
+  79. Justification:
+  80. Guidelines:
+    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
+    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.